### v1

# LAMBench: A Benchmark for Large Atomic Models

Abstract
Large atomic models (LAMs) have undergone remarkable progress recently, emerging as universal or fundamental representations of the potential energy surface defined by the first-principles calculations of atomic systems. However, our understanding of the extent to which these models achieve true universality, as well as their comparative performance across different models, remains limited. This gap is largely due to the lack of comprehensive benchmarks capable of evaluating the effectiveness of LAMs as approximations to the universal potential energy surface. In this study, we introduce LAMBench, a benchmarking system designed to evaluate LAMs in terms of their generalizability, adaptability, and applicability. These attributes are crucial for deploying LAMs as ready-to-use tools across a diverse array of scientific discovery contexts. We benchmark eight state-of-the-art LAMs released prior to April 1, 2025, using LAMBench. Our findings reveal a significant gap between the current LAMs and the ideal universal potential energy surface. They also highlight the need for incorporating cross-domain training data, supporting multi-fidelity modeling, and ensuring the models’ conservativeness and differentiability. As a dynamic and extensible platform, LAMBench is intended to continuously evolve, thereby facilitating the development of robust and generalizable LAMs capable of significantly advancing scientific research. The LAMBench code is open-sourced at https://github.com/deepmodeling/lambench, and an interactive leaderboard is available at https://www.aissquare.com/openlam?tab=Benchmark.
I Introduction
The widespread adoption of large language models (LLMs) is largely driven by the development of general-purpose foundation models pretrained on vast and diverse corpora covering a wide range of disciplines and topics[1]. These foundation models are feasible because there exist common patterns to learn — namely, the shared logic of human language — despite its apparent diversity. In the field of molecular modeling, the fundamental physical principles of quantum mechanics, particularly the Schrödinger equation[2], apply universally to all atomic systems, assuming that relativistic effects are negligible. Under the Born-Oppenheimer approximation[3] a universal potential energy surface (PES) is defined as the ground state solution of the electronic Schrödinger equation, with the nuclear positions treated as input parameters. Consequently, it is feasible to develop a foundational machine learning model to approximate this universal PES. We refer to these molecular foundation models as Large Atomic Models (LAMs) to emphasize their role in capturing fundamental atomic and molecular interactions across diverse chemical systems[4]. LAMs are typically developed through a two-stage process: an initial pretraining phase on broad, diverse atomic datasets to learn a latent representation of the universal PES, followed by fine-tuning on specific downstream datasets to specialize the model for particular target applications.
Despite the existence of a universal solution to the electronic Schrödinger equation, solving it remains computationally demanding even with modern quantum Monte Carlo methods[5]. In practice, Kohn-Sham density functional theory (DFT)[6][7] is the most widely employed computational method for approximating the Born-Oppenheimer PES. The accuracy of DFT calculations is heavily contingent upon the modeling of the exchange-correlation functional, which varies across different research domains. For instance, in materials science, the PBE/PBE+U[8] generalized gradient approximation (GGA) functionals are typically adequate, whereas in chemical science, GGA functionals often fall short, necessitating the use of hybrid functionals[9] for improved accuracy[10]. The disparity in exchange-correlation functionals, along with variations in the choice of basis sets and pseudopotentials, prevents the merging of DFT data across different research domains, thereby impeding the training of a universal potential model.
Nevertheless, domain-specific LAMs are advancing rapidly. For example, MACE-MP-0[11] and SevenNet-0[12] are both trained on the MPtrj dataset[13] from the Inorganic Materials domain at the PBE/PBE+U level of theory. AIMNet[14] and Nutmeg[15] are trained in the domain of small molecules at the SMD(Water)-$\omega$B97X/def2-TZVPP and the $\omega$B97M- D3(BJ)/def2-TZVPPD level of theory, respectively. The rapid advancement of these domain-specific LAMs has transformed the field of atomistic modeling, offering powerful tools for understanding complex inorganic materials and bio-molecular systems. To fully harness the diverse training data from various research domains and maximize the potential of LAMs in learning universal PES, the multitask pretraining strategy presents a promising approach. This strategy encodes shared knowledge into a unified structure with high representational capacity while integrating domain-specific components into multiple neural networks with relatively lower representational power[16, 4]. However, determining the extent to which multitask-trained LAMs approach a truly universal PES remains a challenging question.
Comprehensive and robust benchmarking has proven to be a fundamental prerequisite for the rapid advancement of large-scale machine learning models. For example, benchmarks such as MMLU-Pro[17] and MATH500[18] have driven the rapid progress of LLMs, while the ImageNet[19] benchmark has spurred the rapid iteration of modern computer vision models. Similarly, the CASP benchmark[20] has played a crucial role in advancing protein structure prediction, ultimately leading to the development of AlphaFold2[21].
In the field of molecular modeling, existing benchmarks exhibit two significant limitations. Firstly, they are intrinsically domain-specific, concentrating on isolated sub-fields rather than encompassing a variety of atomic systems. For instance, datasets such as QM9[22] and MD17[23] are used to benchmark molecular property predictions and molecular dynamics (MD) trajectories of small molecules, respectively. These benchmarks are predominantly employed to assess machine learning models within chemical science. The Matbench Discovery[24] evaluates models in the Inorganic Materials domain based on their ability to predict material stability. The Open Catalyst challenges[25] assess models on predicting adsorption energies and relaxed structures for various adsorbate-catalyst combinations. While these benchmarks have played a crucial role in advancing domain-specific LAMs, their fragmented approach undermines the pursuit towards the universal PES model. Secondly, existing assessment methods often fail to reflect real-world application scenarios, reducing their relevance to scientific discovery and technological innovation. For instance, conventional evaluation metrics based on static test sets may not adequately capture the true performance of a model in tasks requiring physically meaningful energy landscapes[26]. Specifically, non-conservative models - where atomic forces are directly inferred from neural networks rather than obtained from the gradient of the predicted energy[27] — can exhibit high apparent accuracy but struggle in applications demanding strict energy conservation, such as MD simulations[28]. The MLIP-Arena benchmark[29] is a step in the direction toward bridging this gap, emphasizing the practical usability of LAMs in tasks such as MD stability and physical property predictions. However, it places less emphasis on evaluating a model’s capacity to generalize across diverse atomic systems or adapt to tasks beyond its training scope, both of which are essential for assessing the performance of LAMs in real scientific discovery.
To address these limitations, we introduce LAMBench, a comprehensive benchmark system designed to rigorously evaluate LAMs across domains, simulation regimes, and application scenarios. Employing LAMBench, we evaluated the performance of eight prominent LAMs released prior to April 1, 2025, uncovering a significant discrepancy between these models and the universal PES. Our findings suggest that enhancing LAM performance requires simultaneous training with data from a diverse array of research domains. Additionally, supporting multi-fidelity at inference time is essential to satisfy the varying requirements of exchange-correlation functionals across different domains. It is also critical to maintain the model’s conservativeness and differentiability to optimize performance in property prediction tasks and ensure stability in molecular dynamics simulations. We believe that the introduction of LAMBench will significantly expedite the development of LAMs, facilitating the creation of ready-to-use models that enhance the pace of real scientific discovery.
II Results
II.1 The LAMBench system
Figure 1: The schematic plot of the LAMBench benchmark.
The LAMBench system is designed to benchmark diverse Large Atomic Models (LAMs) across multiple tasks within a high-throughput workflow, with automation integral to task calculation, result aggregation, analysis, and visualization, as depicted in Figure 1. The implementation details of LAMBench are elaborated in Section IV.2. Key to the system’s effectiveness are the design of the benchmark tasks and the methodologies employed for result interpretation. These benchmark tasks are developed to assess three fundamental capabilities of an LAM: generalizability, adaptability, and applicability. Generalizability pertains to the accuracy of an LAM when utilized as a universal potential across a diverse range of atomic systems. Adaptability denotes the LAM’s capacity to be fine-tuned for tasks beyond potential energy prediction, with particular emphasis on structure-property relationship tasks. Applicability, on the other hand, concerns the stability and efficiency of deploying LAMs in real-world simulations.
Generalizability refers to the accuracy of an LAM on datasets that are not included in the training set. In-distribution (ID) generalizability specifically pertains to the model’s performance on test datasets generated through random splitting from the training datasets, thereby ensuring that these datasets maintain the same distribution as the training data. Conversely, out-of-distribution (OOD) generalizability assesses the model’s performance on test datasets independently constructed, resulting in a distribution distinct from the training data. Within the LAMBench framework, references to generalizability always imply OOD generalizability.
It is important to note that there remains no consensus on the precise definition of OOD or the criteria by which two distributions are considered different. Some researchers define OOD data as those exploring different configurational spaces[30, 31], while others highlight differences in chemical space as critical[25, 32]. In this study, we adopt a practical approach by considering OOD test datasets as downstream datasets designed to address specific scientific challenges, such as training a machine learning potential model for simulating carbon deposition on metal surfaces[33]. These scenarios are most compatible with the downstream applications of LAMs.
In LAMBench, the generalizability is quantitatively assessed using two types of tasks, the force field task and the property calculation task. The force field task assesses the accuracy of LAMs in predicting energy, force, and the virial tensor (when periodic boundary conditions are applied) using 17 datasets that have been utilized in the literature to train machine learning potentials for addressing scientific challenges. Detailed information on these datasets is provided in LABEL:tab:datasets. These test datasets encompass five distinct scientific domains: Inorganic Materials, Catalysis, Reactions, Small Molecules, and Biomolecules and Supramolecules. Given that energy labels calculated via DFT can vary by an arbitrary constant due to variations in pseudopotential selection and software implementations, LAMs are consistently used to predict the energy difference between the label and a dummy model that estimates potential energy solely based on the chemical formula. In contrast, force and virial predictions are directly obtained from the LAMs, as the dummy model invariably yields null predictions for force and virial.
To evaluate the generalizability of LAMs, it is crucial to establish a comprehensive performance indicator that integrates all the energy, force, and virial errors calculated for each configuration across every dataset and domain. An arithmetic mean of the error data is inappropriate due to the differing units of the quantities involved. To address this challenge, we introduce a dimensionless error metric, $\bar{M}^{m}_{\mathrm{FF}}$, with detailed information available in Section IV.3. The error metric is structured so that a dummy model yields a value of $\bar{M}^{m}_{\mathrm{FF}}=1$, whereas an ideal model that perfectly aligns with DFT labels achieves a metric of $\bar{M}^{m}_{\mathrm{FF}}=0$. A lower value of this error metric signifies enhanced generalizability.
Achieving high accuracy in the force field task does not necessarily ensure enhanced performance in property calculations when utilizing the LAM as the force field[26]. For instance, predicting the bulk and shear moduli necessitates finite difference approximations of the second-order derivatives of the potential energy. Consequently, models lacking smoothness up to the second-order derivatives may not consistently yield accurate predictions, even if they perform well in force field tasks. Given the significant variation in investigated properties across different research domains, LAMBench offers a flexible and modular design (see Section IV.2). This design facilitates the implementation of property calculation benchmark tasks, enabling them to be tailored to the specific requirements of diverse research fields.
In this study, we have adapted the MDR Phonon benchmark[34] to evaluate the performance of LAM models in computing key phonon properties, including maximum phonon frequency ($\omega_{\max}$), entropy ($S$), free energy ($F$), and heat capacity at constant volume ($C_{V}$). Additionally, we have utilized the TorsionNet500 benchmark to assess the capability of LAMs in calculating the torsion energy profile of typical drug-like fragments[35]. The Phonon and TorsionNet500 benchmarks are classified as properties of interest within the Inorganic Materials and Small Molecules domains, respectively. These adaptations serve as prototypes for evaluating downstream generalizability on domain-specific property calculations. Further tasks targeting Reactions, Catalysis and other research domains are under active development to align with the evolving capabilities of LAMs. Similar to the approach taken in the force field task, we have introduced a dimensionless error metric, denoted as $\bar{M}^{m}_{\mathrm{PC}}$. This metric is designed to assess the generalizability of an LAM in the property calculation task. For comprehensive details, please refer to Section IV.3.
Adaptability assesses the ability with which a pre-trained LAM can be fine-tuned for tasks beyond its initial training scope, with a particular focus on establishing structure-property relationships. This is distinct from the property calculation task. In the adaptability task, properties are directly predicted by fine-tuned LAMs, whereas, in the property calculation task, properties are computed using LAMs functioning as a force field. The adaption provides a promising strategy for enhancing model accuracy in property prediction tasks, especially in situations where limited training data restricts the achievement of high accuracy through training from scratch[16]. LAMs are typically pre-trained on force field prediction tasks[4, 11, 12] or on denoising tasks[27], which aim to recover stable configurations (local minima of the potential energy surface) from random perturbations in coordinates and atom type masking. Consequently, the ability of LAMs to adapt to property prediction tasks is not straightforward, highlighting the need for a benchmark to evaluate this aspect of LAMs.
Most LAMs comprise a feature extraction module, also referred to as a descriptor, and a fitting module. The feature extraction module encodes information about the universal PES into a latent space during pretraining, while the fitting module decodes this information to perform force field predictions. To predict a new downstream property, another fitting module is randomly initialized and fine-tuned jointly with the feature extraction module, while the original fitting module is discarded. Currently, adaptability tests for LAMs are exclusively supported for models implemented in the DeePMD-kit package[36], while other implementations are not supported.
In this work, we use eight regression tasks from the MatBench benchmark[37] as representative examples. These tasks include the prediction of formation energies for crystal structures and perovskite cells, computed band gaps, exfoliation energies of two-dimensional materials, maximum phonon frequencies of bulk crystalline materials, dielectric constants, and shear moduli. Each task is evaluated using five-fold cross-validation. All properties are treated as intensive quantities, with mean pooling applied to atom-wise predictions. Other tasks, including small molecule property predictions, spectroscopic property predictions, and classification tasks, can be readily integrated into the LAMBench framework in the future.
Finally, applicability assesses the readiness of LAMs for real-world deployment, focusing on computational efficiency and stability. Efficiency typically refers to the time required to compute energy, force, and virial (when applicable) using an LAM on a specific type of computational hardware. Stability examines whether the total energy of a system remains bounded, rather than diverging, during long-timescale MD simulations.
Quantifying the inference efficiency of LAMs in a rigorous and meaningful manner requires careful design, as the observed efficiency can be highly structure-dependent, and different LAMs may exhibit varying performance on the same input. LAMs are designed to be applicable to tasks involving systems of varying sizes, from evaluating the conformational energy of molecules composed of dozens of atoms to large-scale simulations of viruses involving millions of atoms[38]. The efficiency of LAMs—measured as the computational time consumed per atom while determining the energy, forces, and virial of atomic systems—is significantly influenced by the system size and the extent to which the many-core parallelism of modern GPUs is utilized. For relatively small systems, atom-wise efficiency is often reduced due to limited opportunities for parallelization, even though they require less overall computational time for evaluation. As system size increases to approximately 1000 atoms, the average inference time per atom tends to stabilize, as illustrated by Supporting Information Figure S-3. For substantially larger systems, exceeding the memory capacity of a single GPU, multi-GPU parallelism becomes essential. However, such systems are not ideal benchmarks for assessing the efficiency of LAMs, as they are generally computationally intensive, and multi-GPU parallelism is not supported by all the LAMs.
To address the significant size-dependency issues when measuring efficiency in small systems, we concentrate on atomic systems that are large enough to achieve converged efficiency yet small enough to be computed without the need for multi-GPU parallelism. To this end, we randomly sample 1,600 structures from the Inorganic Materials and Catalysis domains in the aforementioned force field task test sets. Each structure is subjected to periodic boundary conditions and duplicated to contain between 800 and 1,000 atoms. This range is sufficiently large to approach the convergence regime while remaining within the memory constraints of most models, thereby avoiding out-of-memory (OOM) errors. Other domains, such as Small Molecules, are excluded from this evaluation, as non-periodic systems generally do not pose sufficient computational demands to effectively stress the inference capabilities of LAMs.
To facilitate the comparison of efficiency across models, we propose a dimensionless efficiency metric, denoted as $M_{\mathrm{E}}^{m}$. This metric is defined as the normalized inverse of the averaged inference time measured for the 1,600 structures. The detailed definition is provided in Section IV.4. A value of $M_{\mathrm{E}}^{m}=1$ corresponds to an efficiency equivalent to a reference value of 0.01 $(\mathrm{\mu s/atom})^{-1}$, with higher value indicating greater efficiency. It is crucial to note that the calculator is fully warmed up by processing 400 structures prior to recording the inference time for the 1,600 structures.
We assessed the stability of LAMs by examining energy conservation in NVE ensemble (microcanonical ensemble) MD simulations across nine atomic systems. These systems were randomly selected from diverse domains, including four periodic structures from the Materials Project[39], three molecular systems from the SPICE2 dataset[40], and two catalytic surface systems from the OC2M dataset[41], as detailed in Supporting Information Figure S-1. For each system, a 10 ps NVE-MD simulation was conducted with a timestep of 1 fs. The initial atomic velocities were sampled from the Boltzmann ensemble at 300 K. The drift of total energy along MD trajectories was quantified using the slope derived from a linear regression applied to the total energy history, with the initial 2 ps of simulation excluded as a warm-up period. Stability is assessed via the instability metric, $M^{m}_{\mathrm{IS}}$, which is calculated based on total energy drift, where smaller values denote improved stability. A detailed definition is provided in Section IV.4.
Table 1: Summary of datasets utilized in force field generalizability assessments. The table details the domain of each test dataset, available labels (E for energy, F for force, and V for virial), number of configurations (frames), exchange-correlation (XC) functional applied for labeling, and additional descriptions. Details on data cleaning procedures can be found in Supporting Information Section S-3.
Name Domain Labels Frames Level of XC Description
Lopanitsyna2023Modeling_A[43] Inorganic Materials EFV 450 PBE-sol Commonly known as HEA25S. A dataset of high-entropy alloy surfaces, focusing on d-block transition metals covering 25 elements.
Lopanitsyna2023Modeling_B[44] Inorganic Materials EFV 25628 PBE-sol Commonly known as HEA25. A dataset of high-entropy alloy bulk structures, focusing on d-block transition metals covering 25 elements.
Dai2024Deep[45] Inorganic Materials EFV 2842 PBE A dataset of high-entropy transition metal diboride (HEMB2) and carbide (HEMC) ceramics.
Torres2019Analysis[46] Inorganic Materials EF 4378 PBE A dataset of Ca-bearing minerals, focusing on silicates and carbonates.
Sours2023Applications[47] Inorganic Materials EF 17414 PBE-D3(BJ) Dataset consisting of 219 different pure silica zeolite topologies.
Batzner2022equivariant[48] Inorganic Materials EF 7499 PBE Down-sampled dataset containing lithium phosphate amorphous glass used by NequIP graph neural network models as evaluation configurations.
WBM_25k[49] Inorganic Materials E 25696 PBE A subset of the WBM dataset randomly down-sampled to 25,696 frames.
Subalex_9k[50] Inorganic Materials EFV 9157 PBE A subset of sAlexandria dataset, randomly down-sampled to 9,157 frames.
ANI-1x[51] Small Molecules EF 8861 $\omega$B97X / 6-31G* A down-sampled dataset from the training data of the ANI-1x potential, containing organic molecule structures.
Zhang2024Active[33] Catalysis EF 452 PBE-D3 Dynamic simulations of carbon deposition on metal surfaces.
Zhang2019Bridging[52] Catalysis EF 9842 PBE Interaction of carbon dioxide with a movable Ni(100) surface.
Vandermause2022Active[53] Catalysis EFV 250 PBE Direct simulation of hydrogen turnover on Pt(111) catalyst surfaces.
Villanueva2024Water[54] Catalysis EF 14859 PBE-D3 Selective CO2 hydrogenation to methanol over oxide catalysts.
Guan2022Benchmark[55] Reactions EF 17412 $\omega$B97X-V / cc-pVTZ Dataset of hydrogen combustion reactions covering 19 reaction channels.
Gasteiger2020Fast[56] Reactions EF 9480 revPBE-D3 / def2-TZVP Validation set from molecular collision experiments with small organic molecules.
MD22[57] Biomolecules EF 2122 PBE+MBD / tighta Molecular dynamics (MD) trajectories of 42-atom tetrapeptide Ac-Ala3-NHMe.
AIMD-Chig[58] Biomolecules EF 19800 M06-2X / 6–31G* A down-sampled MD dataset containing conformations of chignolin protein.
a The “tight” setting corresponds to the default high-accuracy mode in FHI-aims.[42]
II.2 Benchmark results
Table 2 presents the LAMBench Leaderboard, showcasing the performance of LAMs in terms of their generalizability and applicability. Currently, adaptability is excluded from the comparison due to the absence of this test for most of the LAMs under evaluation. This leaderboard aims to strike a balanced comparison between key performance aspects and is intended to offer insights into model suitability for applications relevant to real-world scientific discovery.
Table 2: The LAMBench Leaderboard. Details regarding the calculation of leaderboard metrics are provided in Section IV. $\bar{M}^{m}_{\mathrm{FF}}$ refers to the generalizability error on force field prediction tasks, while $\bar{M}^{m}_{\mathrm{PC}}$ denotes the generalizability error on domain-specific tasks. $M_{\mathrm{E}}^{m}$ stands for the efficiency metric, and $M^{m}_{\mathrm{IS}}$ refers to the instability metric. Arrows alongside the metrics denote whether a higher or lower value corresponds to better performance.
Model Generalizability Applicability
$\bar{M}^{m}_{\mathrm{FF}}\downarrow$ $\bar{M}^{m}_{\mathrm{PC}}\downarrow$ $M_{\mathrm{E}}^{m}\uparrow$ $M^{m}_{\mathrm{IS}}\downarrow$
DPA-2.4-7M 0.265 0.208 0.614 0.039
GRACE-2L-OAM 0.340 0.262 0.678 0.309
SevenNet-l3i5 0.355 0.240 0.279 0.036
MACE-MPA-0 0.356 0.291 0.291 0.000
Orb-v2 0.356 0.560 1.343 2.649
SevenNet-MF-ompa 0.358 0.300 0.088 0.000
MatterSim-v1-5M 0.389 0.280 0.388 0.000
MACE-MP-0 0.405 0.341 0.291 0.089
Figure 2: Dimensionless error metrics for generalizability tasks across different domains. (a) Dimensionless error metrics for force field tasks across five distinct domains. (b) Property-calculation tasks: the Small Molecules domain is assessed using the TorsionNet500 benchmark, while the Inorganic Materials domain is evaluated using the phonon MDR benchmark. To reassess DPA-2.4-7M with enhanced XC functional alignment, referred to as DPA-2.4-7M-bestXC, we selected task heads trained on Yang2023ab for the ANI-1x and AIMD-Chig datasets. For the TorsionNet500 benchmark, we adopted the task head trained on the SPICE2 dataset. For the Sours2023Applications dataset, we applied a D3(BJ) dispersion correction.
In the force field generalizability task, as illustrated in Table 2, performance is evaluated using the dimensionless error $\bar{M}_{\mathrm{FF}}^{m}$. Notably, DPA-2.4-7M demonstrates substantially greater generalizability compared to other LAMs. Interestingly, aside from MatterSim-v1-5M and MACE-MP-0, most models exhibit only minor differences in generalizability. To further investigate the generalizability of these LAMs across specific domains, we present the domain-wise dimensionless error metric in Figure 2(a). In general, most LAMs exhibit relatively low inference errors in the domains of Inorganic Materials, Reactions, and Small Molecules, while showing inferior performance in the domains of Catalysis and Biomolecules and Supramolecules. In the Reactions domain, DPA-2.4-7M, Orb-v2, and SevenNet-MF-ompa achieve the lowest errors, with MatterSim-v1-5M following closely behind. For Small Molecules, DPA-2.4-7M and Orb-v2 rank first and second, respectively, outperforming all other models in this domain. In the Inorganic Materials domain, the comparative performance of the models shows significant alignment with their rankings in the Matbench Discovery benchmark, accessed on April 1st, 2025. The models are ranked by their highest F1 scores as follows, excluding DPA-2.4-7M: SevenNet-MF-ompa $>$ GRACE-2L-OAM $>$ Orb-v2 $>$ MatterSim-v1-5M $>$ MACE-MPA-0 $>$ SevenNet-l3i5 $>$ MACE-MP-0. This coherence further substantiates the reliability and robustness of both the LAMBench and Matbench Discovery benchmark frameworks[24] in evaluating LAMs for Inorganic Materials force field tasks. Within the Catalysis domain, DPA-2.4-7M again achieves the lowest error, significantly surpassing the other models. In the Biomolecules and Supramolecules domain, the DPA-2.4-7M model achieves the highest generalizability, distinctly outperforming all other LAMs. The remaining models show comparable performances, except for MACE-MP-0 and MACE-MPA-0, which fall noticeably behind.
Table 2 reveals that MACE-MPA-0, characterized by a larger parameter set and trained on an extensive dataset, exhibits a dimensionless error 0.05 lower than MACE-MP-0, suggesting enhanced generalizability. This trend is corroborated by the domain-wise error metrics in Figure 2 (a), which demonstrate superior performance across most domains, with the exception of Catalysis, where a minor decline is noted. The Inorganic Materials domain shows the most significant improvement, with an approximately 30 percent reduction in error. Meanwhile, SevenNet-MF-ompa delivers overall performance on par with its lighter counterpart, SevenNet-l3i5; however, its enhanced accuracy in the Inorganic Materials and Reactions domains comes at the cost of reduced generalizability elsewhere, indicating potential overfitting to the Inorganic Materials domain.
The DPA-2.4-7M model demonstrated the best overall performance across all investigated domains. This success can be attributed to its training on 31 datasets, detailed in Supporting Information LABEL:tab:si-datasets, which span multiple domains using a multitask training strategy[16, 4]. In this approach, a common feature extraction structure, the descriptor, is linked to multiple task heads. Each task head is trained on one of the 31 datasets, while the common descriptor is simultaneously trained on all datasets. Consequently, the descriptor is designed to extract the shared knowledge from all the training datasets. For a fair comparison, the task head trained with the MPtrj dataset, which serves as a subset of the training data for all LAMs, is consistently used to assess the generalization error. Thus, the DPA-2.4-7M model’s exceptional performance in the Small Molecules, Biomolecules and Supramolecules, and Catalysis domains can be attributed to the descriptor, which was trained using diverse datasets from domains such as Small Molecules (including the SPICE2 and Yang2023ab datasets) and Catalysis (including the OC20 and OC22 datasets).
It is important to note that certain test datasets are labeled with an exchange-correlation (XC) functional that differs from the one used for labeling the training datasets, such as MPtrj, for the LAMs. For instance, in the domain of Small Molecules, the ANI test set is labeled by $\omega$B97X/6-31G* functional. To further investigate the impact of XC functional mismatch between the model’s training data and the reference functional used in generalizability tests, we reassessed the generalization error of the DPA-2.4-7M model using task heads that align as closely as possible with the XC functional employed for labeling the test datasets. This alignment resulted in a substantial reduction in generalization error across several domains. Specifically, the domain error metric decreased from 0.19 to 0.09 for Small Molecules and from 0.25 to 0.17 for Biomolecules and Supramolecules (see Figure 2 (a)). These findings suggest that XC functional mismatch significantly contributes to the generalization error of the LAM model, highlighting the necessity of training models with multiple XC functional fidelities[59].
The generalizability of LAMs in property calculation tasks is assessed and summarized in Table 2. Compared to force field tasks, property calculation tasks offer greater discriminative power among models. The DPA-2.4-7M model again achieves the best performance, with the lowest error of 0.208—representing nearly a two-thirds reduction in error relative to the worst-performing model. The domain-wise error metric is shown in Figure 2(b), with the detailed raw measurements available in Supporting Information Tables S-3 and S-4.
In the Inorganic Materials domain, where the MDR phonon benchmark is conducted, the SevenNet-MF-ompa, MatterSim-v1-5M, and GRACE-2L-OAM models demonstrate superior performance, with MACE-MPA-0 following closely. Notably, the conservative models exhibit significantly lower errors compared to the non-conservative Orb-v2 model, which directly predicts force. Accurate phonon predictions necessitate the calculation of the force constant matrix, representing the second-order derivative of a force field. Therefore, conservativeness and smoothness are crucial for achieving precise phonon predictions, consistent with previous findings in Ref.[26]. Additionally, the performance difference between MACE-MP-0 and its more expressive counterpart, MACE-MPA-0, suggests that enhanced generalizability in force field prediction tasks can lead to improved performance in a property prediction task.
In the Small Molecules domain, where the TorsionNet500 benchmark is conducted, the dimensionless error values are notably higher compared to those in the Inorganic Materials domain. The multi-task trained LAM, DPA-2.4-7M, demonstrates the lowest error at 0.35, outperforming other models primarily trained on datasets within the Inorganic Materials domain. Utilizing the task head trained with the SPICE2 dataset, which employs a closely aligned XC as the reference DFT method, reduces the error to 0.19. Despite these improvements, performance remains significantly lower than that of domain-specific models such as MACE-OFF23[60], which achieves a dimensionless error of 0.034 on TorsionNet500 benchmark. The discrepancy between the performance of LAMs and MACE-OFF23 highlights the necessity of focusing on domains beyond Inorganic Materials to enhance generalizability in property prediction tasks. This also indicates that current LAMs remain considerably less accurate than domain-specific models and are still distant from achieving the capabilities of a universal PES model.
The adaptation of the pretrained DPA-2.4-7M model across eight Matbench property regression tasks is detailed in Table 3. The fine-tuned model consistently surpasses the model with the same architecture trained from scratch in terms of accuracy. Despite the improvements achieved by the finetuned DPA-2.4-7M model, a discrepancy remains between its performance and the results showcased on the Matbench leaderboard.
Notably, a property predictor adapted from the pretrained MatterSim[61] achieves performance comparable to leading task-specific models on the Matbench leaderboard. Furthermore, JMP, pretrained on the OC20, OC22, ANI-1x, and Transition-1x datasets[16], significantly surpasses existing models, achieving state-of-the-art accuracy. This underscores the critical importance of pretraining for achieving superior model accuracy in downstream property prediction tasks and reinforces the broader vision of LAMs as versatile, high-performing surrogates capable of addressing a wide range of scientific challenges.
Table 3: Adaptability Test: The accuracy of property fine-tuning of DPA-2.4-7M across eight Matbench regression tasks.
Task (Unit) DPA-2.4-7M From scratch DPA-2.4-7M Finetune11footnotemark: 1 Matbench[37] leaderboard MatterSim[61] JMP[16]
MP $E_{\text{form}}$ (meV/atom) 30.1 24.1 17.0 - 10.1
MP Gap (eV) 0.256 0.206 0.156 0.129 0.091
JDFT2D (meV/atom) 43.68 34.09 33.19 32.76 29.94
Phonons (cm-1) 43.55 34.00 28.76 26.02 20.57
Dielectric (unitless) 0.418 0.300 0.271 0.252 0.249
Log KVRH ($\log_{10}\text{GPA}$) 0.061 0.052 0.049 0.049 0.045
Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.068 0.067 0.061 0.059
Perovskites (eV/unitcell) 0.054 0.032 0.027 - 0.026
11footnotemark: 1
Results were obtained using NVIDIA A800-SXM4-40GB GPUs with the maximum permissible batch size. Performance may degrade on NVIDIA L4-24GB GPUs with smaller batch sizes as illustrated in Table S-5, which are used by default in the workflow for cost management purposes.
High efficiency and robust stability are equally critical as model accuracy, especially in the context of high-throughput simulations. The efficiency metric $M^{m}_{\mathrm{E}}$ is summarized in Table 2. Despite having the most parameters, Orb-v2 demonstrates the highest inference efficiency. This superior performance is likely attributed to its non-conservative design, where force predictions are generated from a separate prediction head rather than being derived from energy gradients.
Additionally, we observe that the efficiency is highly sensitive to the test structure for certain LAMs, as exemplified by the broad distribution of the SevenNet-l3i5 model. In contrast, models such as Orb-v2 and DPA-2.4-7M demonstrate relative insensitivity to structural variations. To further investigate the origin of this behavior, a bilayer sodium 2D structure was selected as a representative cases. We systematically reduced the vacuum spacing by shortening the $c$-axis, and the results are summarized in Supporting Information Figure S-4 (a). For Orb-v2 and DPA-2.4-7M, the converged efficiency remained largely unaffected by changes in vacuum spacing. In contrast, other LAMs experienced a significant drop in efficiency as the vacuum was reduced. Once the vacuum reached a certain threshold, further increases had no impact on efficiency. Reducing the vacuum effectively increases the number of neighbors, suggesting that the average number of neighboring atoms within the cutoff radius is the key factor influencing inference efficiency in these models, as demonstrated in Supporting Information Figure S-4 (b). In contrast, Orb-v2 and DPA-2.4-7M utilize a fixed maximum number of neighbors through padding, rendering them insensitive to such variations.
Figure 3: Distribution of inference time, normalized by the number of atoms, measured across 1,600 randomly selected configurations. Lower values indicate higher efficiency.
The stability of the LAMs is evaluated using the instability metric $M^{m}_{\mathrm{IS}}$, as presented in Table 2. The conservative models, such as MatterSim-v1-5M, exhibit minimal energy drift throughout the evaluation, indicating stable simulations. In contrast, the non-conservative model, Orb-v2, shows instability metric several orders of magnitude larger, reflecting instability in the simulation. In general, models with a greater number of trainable parameters and trained on larger datasets exhibit higher stability, as reflected in reduced energy drift. This trend is evident in the MACE, and SevenNet model families as shown in Supporting Information Table S-6.
III Discussion and Conclusion
This study introduces a comprehensive benchmarking framework, designated as LAMBench, for evaluating large atomic models (LAMs). It is designed to assess the extent to which LAMs can be used as versatile, out-of-the-box tools capable of advancing scientific discovery across a broad range of contexts. We emphasize three essential requirements for realizing this vision: generalizability to atomic systems across diverse research domains, adaptability to novel tasks, and applicability in real-world simulations with regard to stability and efficiency.
We propose two test tasks to evaluate generalizability: the force field tasks, which assess LAMs’ accuracy in predicting energy, force, and virial (when applicable), and the property calculation tasks, which measure accuracy in computing properties of interest for specific applications using the LAM as a force field. The errors of LAMs in these tasks are interpreted using a dimensionless error metric. This metric facilitates the comparison of generalizability across different test tasks, despite significant variations in label magnitudes, and offers a clear indication of the model’s proximity to an ideal universal model.
In this study, we benchmarked eight leading LAMs, released before April 1, 2025. DPA-2.4-7M consistently demonstrated superior performance in terms of generalizability compared to all other LAMs. In the force field generalizability test, DPA-2.4-7M was followed by GRACE-2L-OAM, SevenNet-l3i5, SevenNet-MF-ompa, MACE-MPA-0, and Orb-v2, all exhibiting remarkably similar accuracy levels. A trend of overfitting to the Inorganic Materials domain was observed in some LAMs. However, in the generalizability test for property calculation, more pronounced differences between the models are evident. In the phonon calculation test, the disparities are mainly attributed to the conservativeness of the models, with conservative models significantly outperforming the non-conservative ones. In the torsion profile calculation task within the Small Molecules domain, where most LAMs, except DPA-2.4-7M and SevenNet-l3i5, exhibit a dimensionless error of $\geq 0.5$. Although the XC-adapted DPA-2.4-7M model achieves the lowest generalization error of 0.19, it remains significantly higher than MACE-OFF23, a model specifically trained for the Small Molecules domain, which has an error of 0.034.
The LAMBench results highlight a substantial gap between current LAMs and the ideal universal model, envisioned as an out-of-the-box simulation tool for scientific discovery. This gap is primarily due to limited generalizability beyond the Inorganic Materials domain, which is largely attributed to training datasets predominantly sourced from this domain. While DPA-2.4-7M employs a multi-task training scheme, simultaneously utilizing data from multiple domains to enhance generalization capabilities, its accuracy remains inferior to models specifically oriented towards individual domains. To bridge this gap, future research might focus on developing advanced model architectures for improved generalizability, more efficient training methods to learn shared knowledge across diverse domains, and a balanced distribution of training data across various research fields. These directions hold promise for advancing the capabilities of LAMs in scientific discovery.
The capability to adjust model fidelity is proposed as an essential feature for LAMs, given that different research domains necessitate varying levels of accuracy in DFT calculations. For instance, the Inorganic Materials domain typically requires GGA functionals, whereas the Small Molecules domain often demands hybrid functionals. Notably, the force field error in the Small Molecules domain for DPA-2.4-7M decreases from 0.19 to 0.09 when switching from PBE/PBE+U to the $\omega$B97X-D functional. This underscores the importance of adaptable fidelity in achieving optimal performance across diverse research applications.
Regarding the efficiency of using LAMs in simulations, the non-conservative model Orb-v2 demonstrates significantly higher performance compared to the conservative models. Among the conservative models, the fastest model is 7.7 times quicker than the slowest, a disparity that is notably larger than the variations observed in generalizability. This observation underscores the importance of considering the efficiency in the development of LAMs, particularly for speed-critical applications such as long-time molecular dynamics simulations.
In terms of stability, conservative models exhibit significantly superior performance compared to non-conservative models, a finding that is corroborated by existing literature[26]. This observation, along with the necessity for conservativeness in phonon calculation tasks, emphasizes the importance of integrating conservativeness into the design of LAMs.
In this study, adaptability is exclusively benchmarked for the DPA-v2.4-7M model, as most implementations of LAMs are primarily confined to force field prediction and are not readily adaptable to property prediction tasks. The adaptability test for DPA-2.4-7M demonstrates that a pretrained model can offer substantial advantages in downstream property prediction tasks compared to models trained from scratch, particularly in data-scarce scenarios. Despite DPA-2.4-7M’s strong performance in generalizability tests, a noticeable gap remains between its accuracy and those of state-of-the-art property prediction models. This highlights a significant opportunity for further enhancing the adaptability of the DPA-2.4-7M model representing a crucial step towards the development of ideal universal models. Furthermore, it is advisable for LAM developers to enhance their models with the capability to adapt for property prediction. This advancement could open up new avenues for their application in scientific discovery.
In the future, LAMBench should be enhanced to more accurately reflect the LAMs’ performance in real-world applications as an out-of-box simulation tool. This enhancement requires the collection of additional test datasets for the force field generalizability test and the incorporation of more property calculation workflows in the property generalizability test. It is important to note that the design of property calculation tasks requires careful consideration. Directly borrowing all property calculations from applications may not be optimal, as these calculations often necessitate long time-scale and large spatial scale MD simulations, such as for ionic diffusion constants, which demand substantial computational resources. Ideally, property benchmarks should be designed as representative examples that reflect performance in the calculation of domain-specific properties while remaining computationally feasible.
Evaluating adaptability presents challenges, even when LAM implementations support fine-tuning for property prediction tasks. The adaptability benchmark necessitates a fine-tuning procedure, meaning that comparisons of final performance are influenced not only by the model’s adaptability but also by variations in fine-tuning code. Establishing a unified property fine-tuning framework for all LAMs could be the most effective solution; however, this would require extensive development and is beyond the current scope of LAMBench.
Given the diverse application scenarios of LAMs, LAMBench is designed as a dynamic system, continuously incorporating additional test tasks and datasets over time. Establishing such a framework requires sustained, community-driven efforts and consensus; therefore, we strongly encourage ongoing contributions from the community.
IV Method
IV.1 Models
To evaluate the thoroughness and discriminative power of the LAMBench benchmark, we test a series of LAMs, as listed in Table 4. Most models follow a single-task, single-fidelity training strategy, using datasets curated under consistent DFT settings—primarily from the domain of Inorganic Materials. Exceptions include SevenNet-MF-OMPA and DPA-2.4-7M. SevenNet-MF-OMPA employs multi-fidelity training on the OMat24[50], MPtrj, and sAlex[50] datasets, achieving high accuracy despite heterogeneity in DFT settings, though still focused on inorganic systems. In contrast, DPA-2.4-7M adopts a multitask training approach using the OpenLAM dataset collection (Supporting Information LABEL:tab:si-datasets), which spans a broad range of chemical and material systems—including catalysis, small molecules, reactions, and biomolecules—under diverse DFT settings. Detailed hyperparameter information for the models can be found in Supporting Information Table S-1. The multitask training of DPA-2.4-7M produces 31 prediction heads, each corresponding to a distinct training task within OpenLAM. Unless otherwise noted, the prediction head trained by the MPtrj dataset is used by default.
Table 4: Summary of LAMs benchmarked in this study. The table includes the model name, number of parameters, training dataset, and cutoff radius. For conservative models that calculate force as the negative gradient of energy, “Direct Force Prediction” is indicated as “No”.
Model # Parameters Training Set Direct Force Prediction Cutoff Radius (Å)
DPA-2.4-7M[4] 6.64M OpenLAM No 6.0
MACE-MP-0 medium[11] 4.69M MPtrj No 6.0
MACE-MPA-0 medium[11] 9.06M MPtrj, sAlex No 6.0
Orb-v2[27] 25.2M MPtrj, Alex Yes 10.0
SevenNet-MF-ompa[59] 25.7M OMat24, MPtrj, sAlex No 6.0
SevenNet-l3i5[12] 1.17M MPtrj No 5.0
MatterSim-v1-5M[61] 4.55M MattterSim No 5.0
GRACE-2L-OAM[62] 15.3M MPtrj No 6.0
IV.2 LAMBench Implementation
Benchmarking LAMs involves repeatedly performing computations using various combinations of models and test tasks, followed by aggregating and visualizing the benchmarking results. The combination of models and tasks generates a substantial job array, rendering manual submission inefficient. To automate and enhance the benchmarking process, we developed the LAMBench-toolkit, as illustrated in Figure 1. By offering a structured and automated benchmarking framework, LAMBench significantly facilitates the comprehensive evaluation and comparison of LAMs. The LAMBench-toolkit is openly available under the MIT License at github.com/deepmodeling/lambench. An interactive leaderboard is provided at https://www.aissquare.com/openlam?tab=Benchmark.
In the LAMBench-toolkit, model definitions, test tasks, and workflow management are implemented as distinct modules. This modular design allows LAM developers to effortlessly incorporate new models and test cases into the toolkit, while also facilitating the efficient maintenance of existing components.
Models.
Each LAM within the LAMBench-toolkit is specified via a configuration file that includes the associated Python package name and the path for loading model weights. Models engage with test tasks through the Atomic Simulation Environment (ASE) calculator interface[63], offering a standardized approach for model-task interaction. Developers can seamlessly integrate new LAMs into LAMBench-toolkit by providing their ASE calculators.
Tasks.
The tasks module implements the benchmark tasks that researchers intend to perform on LAMs. Each task explicitly delineates the calculation workflow to evaluate a specific capability of a model and provides output metrics, such as the error in calculating a particular value, to quantify performance. Upon completion of a task, the resulting metrics and model information are uploaded to a database, facilitating easier data analysis and management. Additionally, the database is utilized to identify and skip duplicate computational jobs at the start of each task.
Workflow.
Within the context of benchmarking LAMs, a computational workflow denotes a structured sequence of computational steps aimed at assessing model performance across diverse models and datasets. The workflow module of LAMBench orchestrates these benchmarking steps, efficiently handling job submissions, executions, and the subsequent aggregation and analysis of results. This design ensures that developers of models and test tasks are not burdened by the specifics of execution on computational resources, nor are they required to manually collect and analyze the results.
The array of computational steps generated by the workflow module is submitted to computational resources through Dflow[64], a functional programming interface designed for scientific computing workflows. In our experiments, jobs are executed on cloud instances equipped with an NVIDIA V100 32GB GPU via the Bohrium Cloud Platform. Various other computational resources, including high-performance supercomputers and local hardware, are also supported. Upon job completion, the workflow retrieves results from the database, calculates the metrics, and updates the visualization plots on the webpage frontend. This process enables researchers to intuitively analyze and interpret the benchmarking outcomes.
IV.3 Generalizability metrics
In assessing the generalizability of models, the primary metrics employed are the mean absolute error (MAE) and root mean square error (RMSE) for specific predictions across test sets within various domains. Direct comparison of model performance using these metrics can be challenging due to the extensive number of metrics generated for each model. This complexity arises because each model typically provides numerous error measurements across different prediction types and test sets. Thus, a more integrated approach to evaluation is required to effectively compare the generalizability of different models, considering the multitude of error metrics involved.
[…]
$\hat{M}^{m}_{k,p,i}=\frac{M^{m}_{k,p,i}}{M^{\mathrm{dummy}}_{k,p,i}}$ (1)
This baseline model predicts energy based solely on the chemical formula, disregarding any structural details, thereby providing a reference point for evaluating the improvement offered by more sophisticated models.
For each domain, we compute the log-average of normalized metrics across all datasets within this domain by
$\bar{M}^{m}_{k,p}=\exp\left(\frac{1}{n_{k,p}}\sum_{i=1}^{n_{k,p}}\ln\hat{M}^{m}_{k,p,i}\right),$ (2)
where $n_{k,p}$ denotes the number of test sets for domain $k$ and prediction type $p$. Subsequently, we calculate a weighted dimensionless domain error metric to encapsulate the overall error across various prediction types:
$\bar{M}^{m}_{k}=\sum_{p}w_{p}\bar{M}^{m}_{k,p}\Bigg{/}\sum_{p}w_{p},$ (3)
where $w_{p}$ denotes the weights assigned to each prediction type $p$.
Finally the overall generalizability error metric of a model across all the domains is defined by the average of the domain error metrics,
${\bar{M}^{m}}=\frac{1}{n_{D}}\sum_{k=1}^{n_{D}}{\bar{M}^{m}_{k}},$ (4)
where $n_{D}$ denotes the number of domains under consideration. The dimensionless error metric ${\bar{M}^{m}}$ allows for the comparison of generalizability across different models. It reflects the overall generalization capability across all domains, prediction types, and test sets, with a lower value indicating superior performance. The only tunable parameter is the weights assigned to prediction types, thereby minimizing arbitrariness in the comparison system.
For the force field generalizability tasks, we adopt RMSE as the primary error metric. The prediction types include energy and force, with weights assigned as $w_{E}=w_{F}=0.5$. When periodic boundary conditions are assumed and virial labels are available, virial predictions are also considered. In this scenario, the prediction weights are adjusted to $w_{E}=w_{F}=0.45$ and $w_{V}=0.1$. The resulting error metric is referred to as $\bar{M}^{m}_{\mathrm{FF}}$.
For the domain-specific property calculation tasks, we adopt the MAE as the primary error metric. In the Inorganic Materials domain, the MDR phonon benchmark predicts maximum phonon frequency, entropy, free energy, and heat capacity at constant volume, with each prediction type assigned a weight of 0.25. In the Small Moleculesdomain, the TorsionNet 500 benchmark predicts the torsion profile energy, torsional barrier height, and the number of molecules for which the model’s prediction of the torsional barrier height has an error exceeding 1 kcal/mol. Each prediction type in this domain is assigned a weight of $\frac{1}{3}$. The resulting error metric is denoted as $\bar{M}^{m}_{\mathrm{PC}}$.
IV.4 Applicability metrics
The applicability metrics incorporate both efficiency and stability and are computed differently from the generalizability metrics due to the absence of a dummy model baseline. To evaluate efficiency, we define an efficiency metric, $M_{E}^{m}$ , by normalizing the average inference time (with unit $\mathrm{\mu s/atom}$), $\bar{\eta}^{m}$, of a given LAM measured over 1600 configurations with respect to an artificial reference value, thereby rescaling it to a range between zero and positive infinity. A larger value indicates the higher efficiency.
$M_{\mathrm{E}}^{m}=\frac{\eta^{0}}{\bar{\eta}^{m}},\quad\eta^{0}=100\ \mathrm{\mu s/atom},\quad\bar{\eta}^{m}=\frac{1}{1600}\sum_{i}^{1600}\eta_{i}^{m},$ (5)
where $\eta_{i}^{m}$ is the inference time of configuration $i$ for model $m$.
Stability is quantified by measuring the total energy drift in NVE simulations across nine structures. For each simulation trajectory, an instability metric is defined based on the magnitude of the slope obtained via linear regression of total energy per atom versus simulation time. A tolerance value, $5\times 10^{-4}\ \mathrm{eV/atom/ps}$, is determined as three times the statistical uncertainty in calculating the slope from a 10 ps NVE-MD trajectory using the MACE-MPA-0 model. If the measured slope is smaller than the tolerance value, the energy drift is considered negligible. We define the dimensionless measure of instability for structure $i$ as follows:
$M^{m}_{\mathrm{IS},i}=\begin{cases}\max\left(0,\log_{10}({\Phi_{i}}/{\Phi_{\mathrm{tol}}})\right),&\text{if success}\\ 5,&\text{otherwise}\end{cases},\quad\Phi_{\mathrm{tol}}=5\times 10^{-4}\ \mathrm{eV/atom/ps},$ (6)
where $\Phi_{i}$ represents the total energy drift, and $\Phi_{\mathrm{tol}}$ denotes the tolerance. This metric indicates the relative order of magnitude of the slope compared to the tolerance. In cases where a MD simulation fails, a penalty of 5 is assigned, representing a drift five orders of magnitude larger than the typical statistical uncertainty in measuring the slope. The final instability metric is computed as the average over all nine structures.
$M^{m}_{\mathrm{IS}}=\frac{1}{9}\sum_{i=1}^{9}M^{m}_{\mathrm{IS},i}$ (7)
This result is bounded within the range $[0,+\infty)$, where a lower value signifies greater stability.
Acknowledgments
The computational resources utilized in this work were provided by the AI for Science Institute. The work is supported by National Key Research and Development Program of China Grant No. 2022YFA1004300. A.L. acknowledges funding from the Horizon Europe MSCA Doctoral network grant n.101073486, EUSpecLab, funded by the European Union.
Supporting Information
S-1 DPA Model Configurations
Table S-1: Hyperparameters for DPA-2.4-7M model
Module Hyperparameter DPA-2.4-7M Repinit nsel 120 neuron 3$\times$64 three_body_neuron 3$\times$32 Repformer nsel 40 nlayers 6 g1_dim 384 g2_dim 96 Activation tanh Training batch size “auto:256” nGPUs 120 $\times$ A800-40GB start lr 1e-3 stop lr 1e-5 precision float32 steps trained 8M
Table S-2: Summary of datasets including dataset name, frame numbers, data cleaning, training average atoms, and DFT level.
Dataset Name Frames Train Avg Atoms Data Cleaning Level of Theory
Alex2D[65] trn: 1,223,831 val: 135,810 9.9 Exclude frames with energies $>$ 0 eV/atom, maximum absolute force $>$ 5 eV/Å, maximum absolute virial/atom $>$ 8 eV/atom; Trajectory first/last 3 frames + every 10 frame PBE/PAW, 520 eV, 0.4 Å-1
MPTraj[13] trn: 1,401,956 val: 110,918 31.35 According to the distribution difference between MACE-MP-0 prediction and label, the union set of top 0.1% with the largest uncorrected energy difference and the top 0.1% with the largest force difference were excluded, resulting in a total of 3123 frames excluded. PBE (+U)/PAW, 520 eV, 0.04 Å-1
OC20M[66] trn: 20,000,000 val: 999,866 73.1 - rPBE/PAW, 350eV
OC22[67] trn: 8,194,770 val: 394,727 79.8 Exclude frames with energies $>$ 0eV/atom, maximum absolute force $>$ 10eV/Å PBE (+U)/PAW, 350 eV, 0.157 Å-1
ODAC23[68] trn: 2,682,332 val: 63,623 203 Exclude frames with energy/atom $>$ 0.5 eV/atom or $<$ -0.2 eV/atom, and maximum absolute force $>$ 25 eV/Å; Trajectory first/last 3 frames + every 20th frame PBE-D3/PAW, 600 eV, gamma
OMAT24[50] trn: 100,568,820 val: 1,074,647 18.6 Exclude frames with energies $>$ 0 eV/atom, $<$ -25eV/atom, maximum absolute force $>$ 50eV/Å PBE (+U)/PAW
SPICE2[69] trn: 1,621,168 val: 180,089 35.6 Exclude frames with energies $<$ -10000 eV/atom, maximum absolute force $>$ 15eV/Å $\omega$B97M-D3(BJ)/def2-TZVPPD
Transition1x[70] trn: 7,632,328 val: 967,454 13.91 - $\omega$B97X/6-31G(d)
UniPero[71] trn: 14,487 val: 1,357 53.64 - PBEsol/LCAO, 1360 eV, 0.189 Å-1
Dai2023Alloy[72] trn: 71,482 val: 1,240 20.99 - PBE/Norm-conserving, 1360 eV, 0.094 Å-1
Zhang2023Cathode[73] trn: 88,692 val: 9,695 46.28 - PBE (+U)/PAW, 520 eV, 0.25 Å-1
Gong2023Cluster[74] trn: 143,418 val: 15,331 23.67 - PBE-D3/TZV2P, 400-1000 Ry
Li2025APEX[75] trn: 24,097 val: 100 24.03 - PBE/NC, 1360 eV, 0.15 Å-1
Shi2024Electrolyte[76] trn: 65,393 val: 3,438 192.69 - PBE-D3, 800 Ry
Shi2024SSE[77] trn: 125,083 val: 6,587 48.33 - PBE-sol/LCAO, 1360 eV, 0.28 Å-1
Yang2023ab[78] trn: 1,379,956 val: 24,257 33.47 - $\omega$B97X-D/6-31G**
Li2025General[79] trn: 14,024,587 val: 1,558,260 17.8 Exclude maximum absolute force $>$ 20eV/Å GFN2-xTB
Huang2021Deep-PBE[80] trn: 17,582 val: 886 218.59 - PBE/PAW, 650 eV, 0.26 Å-1
Liu2024Machine[81] trn: 215,481 val: 23,343 70.36 - PBE/LCAO, 1360 eV, 0.151 Å-1
Zhang2021Phase[82] trn: 46,077 val: 2,342 172.3 - SCAN/PAW, 1500 eV, 0.5 Å-1
Jiang2021Accurate[83] trn: 138,194 val: 3,965 23.16 - PBE/PAW, 650 eV, 0.1 Å-1
Chen2023Modeling[84] trn: 6,449 val: 276 16.94 - SCAN/PAW, 650 eV, 0.1 Å-1
Unke2019PhysNet[85] trn: 2,594,609 val: 136,571 21.38 - revPBE-D3(BJ)/def2-TZVP
Wen2021Specialising[86] trn: 10,054 val: 474 20.02 - PBE/PAW, 650 eV, 0.1 Å-1
Wang2022Classical[87] trn: 14,935 val: 738 24.47 - PBE/PAW, 650 eV, 0.1 Å-1
Wang2022Tungsten[88] trn: 42,297 val: 2,100 24.91 - PBE/PAW, 600 eV, 0.16 Å-1
Wu2021Deep[89] trn: 27,660 val: 917 96 - PBE/PAW, 600 eV, 2$\times$2$\times$2 kpt
Huang2021Deep-PBEsol[80] trn: 7,502 val: 384 160.8 - PBE-sol/PAW, 650 eV, 0.26 Å-1
Wang2021Generalizable[90] trn: 64,239 val: 2,256 21.85 - PBE-D3/PAW, 650 eV, 0.1 Å-1
Wu2021Accurate[91] trn: 11,621 val: 568 21.73 - PBE/PAW, 700 eV
Tuo2023Hybrid[92] trn: 48,078 val: 2,530 45.11 - PBE-D3(BJ)/PAW, 500 eV, 0.16 Å-1
S-2 Original Error Metrics for Generalizability Domain Specific Tasks
Table S-3: Phonon related property prediction on MDR phonon benchmark.
Model MAE $\omega_{\text{max}}$ (K) MAE $S$ (J/K/mol) MAE $F$ (kJ/mol) MAE $C_{V}$ (J/K/mol)
DPA-2.4-7M 58.9 45.6 18.1 12.8
MACE-MP-0 61.0 59.6 23.8 13.1
MACE-MPA-0 29.7 19.8 7.9 5.8
Orb-v2 308.0 446.5 184.3 58.5
SevenNet-l3i5 25.6 25.9 9.6 4.9
SevenNet-MF-ompa 14.9 10.5 4.1 3.1
MatterSim-v1-5M 16.4 15.2 5.2 3.1
GRACE-2L-OAM 19.5 14.1 5.5 3.7
dummy 1188.3 764.8 125.1 547.4
Table S-4: Torsional MAE and MAEB errors between LAM predictions and its reference DFT labels on the TorsionNet-500 Benchmark.
Model MAE (kcal/mol) MAEB11footnotemark: 1 (kcal/mol) $\text{NABH}_{h}$22footnotemark: 2
DPA-2.4-7M 0.727 1.220 267
MACE-MP-0 1.653 2.425 354
MACE-MPA-0 1.486 2.179 339
Orb-v2 1.238 2.067 345
SevenNet-l3i5 1.113 1.641 300
SevenNet-MF-ompa 1.599 2.309 354
MatterSim-v1-5M 1.348 2.240 345
GRACE-2L-OAM 1.387 1.992 303
dummy 2.501 5.877 494
11footnotemark: 1
The mean-absolute-error of the torsional barrier height, defined as the difference between the minimum and the maximum energy points during the torsional rotation. 22footnotemark: 2The number of molecules (total: $N_{\text{mols}}$ = 500) for which the model prediction of torsional barrier height has an error of more than 1 kcal/mol.
S-3 Downsampling Details for OOD Datasets
We applied a variety of down-sampling strategies across eleven benchmark datasets to reduce computational cost while preserving representativeness.
The datasets employed in this study underwent various downsampling procedures to ensure manageable sizes while maintaining representative distributions. For the ANI-1x dataset, 10% of frames were randomly sampled from the validation set using a fixed random seed 42. The Lopanitsyna2023Modeling_B dataset was used in its entirety without downsampling, while its surface counterpart Lopanitsyna2023Modeling_A utilized the validation split provided by the original authors. The combined Dai2024Deep dataset was reduced to 5% of its original size through random sampling with seed 0. For the MD22 dataset, 5% of frames were randomly sampled from systems of short peptides (Ac-Ala3-NHMe), docosahexaenoic acid, stachyose, and DNA base pairs (AT-AT and AT-AT-CG-CG), with random seed set to 0 and units fixed. The Gasteiger2020Fast dataset retained its original test split without any downsampling. The Guan2022Benchmark dataset was 5% subset randomly sampled (seed 0) from the pre-cleaned data which excluding frames with energy $>-0.5~{}\mathrm{eV/atom}$ and maximum force $>30~{}\mathrm{eV/\text{\AA}}$. The Zhang2019Bridging dataset excluded frames with energy $<0.53~{}\mathrm{eV/atom}$ and maximum forces $>5~{}\mathrm{eV/\text{\AA}}$. The Batzner2022equivariant dataset was downsampled to 5% (seed 0) after removing frames with energy $<-50~{}\mathrm{eV/atom}$ and maximum forces $>10~{}\mathrm{eV/\text{\AA}}$. The AIMD-Chig dataset underwent downsampling at 0.5% (seed 0). The Zhang2024Active dataset was randomly downsampled to 5% (seed 0) after removing frames with energy $>-2.5~{}\mathrm{eV/atom}$ and maximum forces $>25~{}\mathrm{eV/\text{\AA}}$. The Villanueva2024Water dataset was used in full after removing frames with energy ($<-6~{}\mathrm{eV/atom}$) and force ($<2~{}\mathrm{eV/\text{\AA}}$). The Subalex_9k dataset was created by randomly sampling 9157 frames (seed 42) from pre-cleaned data that excluded frames with energy $<-25~{}\mathrm{eV/atom}$ and maximum forces $>3~{}\mathrm{eV/\text{\AA}}$ from sAlex. The WBM_25k dataset was downsampled to 25696 frames (seed 42). The Sours2023Applications dataset was reduced to 5% (seed 0) after removing frames with energy $>-5~{}\mathrm{eV/atom}$. The Torres2019Analysis (Ca_Battery) was used in all GGA-labeled dataset with no virials. The Vandermause2022Active was used in its full set of DFT labels.
S-4 Structures for Conservativeness Benchmark
Figure S-1: Structures used for the Conservativeness Benchmark. (a) H2Al32Cr48Mn16N2O; (b) Cs8N2; (c) Gd2Ni2Si4; (d) NdPd3; (e) BaNi2O8V2; (f) BaNiO5Tm2; (g) CH3N5S; (h) C3H5N2; (i) C4H7NO.
S-5 Inference Efficiency Convergence Test
Figure S-2: The structures in unit cell for convergence test in efficiency benchmark. (a) HCuMg11O12; (b) Na; (c) High-entropy alloy: Ag2AuCoCrCuFe2HfIrLuMnMo2Nb2Ni2PdPt2RhRuSc2Ta2Ti2VW3YZnZr; (d) H2ClCr2FO10Pb4; (e) BN; (f) CNi36O2.
Figure S-3: The convergence test of efficiency benchmark with respect to atom numbers.
S-6 The Coordination Number Analysis of Configurations in Efficiency Benchmarks
Figure S-4: Effect of the number of neighbor atoms on inference time, (a) correlation between inference time and vacuum thickness in a bilayer $\text{Na}_{8}$ 2D structure. (b) Dependence of inference time on the number of neighbor atoms in a bilayer $\text{Na}_{8}$ 2D structure. The number of neighbor atoms is estimated using the cutoff radius specific to each model.
S-7 Property Fine-tuning Batch Size Effect
As mentioned earlier, prediction accuracy is strongly correlated with the training batch size due to the limited size of the fine-tuning dataset. The default workflow utilizes NVIDIA L4 GPUs for cost management purposes, which results in a notable decline in performance. However, the relative performance trends among models remain consistent. For the MP $E_{\text{form}}$ and MP Gap tasks, the reported accuracy may not reflect the fully converged results due to insufficient training epochs.
Table S-5: Comparison of property fine-tuning accuracy on NVIDIA L4-24GB and A800-40GB GPU. For MP $E_{\text{form}}$ and MP Gap, the batch size on the L4 was set to a quarter of that used on the A100, while for all other tasks, it was set to half.
Tasks (Units) DPA-2.4-7M L4 Finetune DPA-2.4-7M A800 Finetune Estimated Epochs11footnotemark: 1 L4 |||| A800 MP $E_{\text{form}}$ (meV/atom) 31.1 24.1 80 |||| 320 MP Gap (eV) 0.285 0.206 100 |||| 400 JDFT2D (meV/atom) 46.93 34.09 28500 |||| 57000 Phonons (cm-1) 41.10 34.00 14000 |||| 28000 Dielectric (unitless) 0.431 0.300 3000 |||| 6000 Log KVRH ($\log_{10}\text{GPA}$) 0.062 0.052 1300 |||| 2600 Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.068 1300 |||| 2600 Perovskites (eV/unitcell) 0.061 0.032 1350 |||| 2700
11footnotemark: 1
The DPA models employ a dynamic batch size to optimize training efficiency, making it difficult to precisely determine the number of training epochs.
S-8 Stability Test Results
Table S-6: NVE molecular dynamics simulations over nine atomic systems.
Model Energy Drift (meV/atom/ps) Success Rate Direct Force Prediction
DPA-2.4-7M 0.018 1 No
MACE-MP-0 0.005 1 No
MACE-MPA-0 0.004 1 No
Orb-v2 222.8 1 Yes
SevenNet-l3i5 0.012 1 No
SevenNet-MF-ompa 0.010 1 No
MatterSim-v1-5M 0.007 1 No
GRACE-2L-OAM 0.010 1 No

### v2

@@
-LAMBench: A Benchmark for Large Atomic Models
+LAMBench: A Benchmark for Large Atomistic Models
 Abstract
-Large atomic models (LAMs) have undergone remarkable progress recently, emerging as universal or fundamental representations of the potential energy surface defined by the first-principles calculations of atomic systems.
+Large Atomistic Models (LAMs) have undergone remarkable progress recently, emerging as universal or fundamental representations of the potential energy surface defined by the first-principles calculations of atomistic systems.
 However, our understanding of the extent to which these models achieve true universality, as well as their comparative performance across different models, remains limited.
@@
 These attributes are crucial for deploying LAMs as ready-to-use tools across a diverse array of scientific discovery contexts.
-We benchmark eight state-of-the-art LAMs released prior to April 1, 2025, using LAMBench.
+We benchmark ten state-of-the-art LAMs released prior to August 1, 2025, using LAMBench.
 Our findings reveal a significant gap between the current LAMs and the ideal universal potential energy surface.
@@
 These foundation models are feasible because there exist common patterns to learn — namely, the shared logic of human language — despite its apparent diversity.
-In the field of molecular modeling, the fundamental physical principles of quantum mechanics, particularly the Schrödinger equation[2], apply universally to all atomic systems, assuming that relativistic effects are negligible.
+In the field of molecular modeling, the fundamental physical principles of quantum mechanics, particularly the Schrödinger equation[2], apply universally to all atomistic systems, assuming that relativistic effects are negligible.
 Under the Born-Oppenheimer approximation[3] a universal potential energy surface (PES) is defined as the ground state solution of the electronic Schrödinger equation, with the nuclear positions treated as input parameters.
 Consequently, it is feasible to develop a foundational machine learning model to approximate this universal PES.
-We refer to these molecular foundation models as Large Atomic Models (LAMs) to emphasize their role in capturing fundamental atomic and molecular interactions across diverse chemical systems[4].
+We refer to these molecular foundation models as Large Atomistic Models (LAMs) to emphasize their role in capturing fundamental atomic and molecular interactions across diverse chemical systems[4].
 LAMs are typically developed through a two-stage process: an initial pretraining phase on broad, diverse atomic datasets to learn a latent representation of the universal PES, followed by fine-tuning on specific downstream datasets to specialize the model for particular target applications.
@@
 The disparity in exchange-correlation functionals, along with variations in the choice of basis sets and pseudopotentials, prevents the merging of DFT data across different research domains, thereby impeding the training of a universal potential model.
-Nevertheless, domain-specific LAMs are advancing rapidly.
-For example, MACE-MP-0[11] and SevenNet-0[12] are both trained on the MPtrj dataset[13] from the Inorganic Materials domain at the PBE/PBE+U level of theory.
-AIMNet[14] and Nutmeg[15] are trained in the domain of small molecules at the SMD(Water)-$\omega$B97X/def2-TZVPP and the $\omega$B97M- D3(BJ)/def2-TZVPPD level of theory, respectively.
-The rapid advancement of these domain-specific LAMs has transformed the field of atomistic modeling, offering powerful tools for understanding complex inorganic materials and bio-molecular systems.
-To fully harness the diverse training data from various research domains and maximize the potential of LAMs in learning universal PES, the multitask pretraining strategy presents a promising approach.
-This strategy encodes shared knowledge into a unified structure with high representational capacity while integrating domain-specific components into multiple neural networks with relatively lower representational power[16, 4].
-However, determining the extent to which multitask-trained LAMs approach a truly universal PES remains a challenging question.
+[…]
 Comprehensive and robust benchmarking has proven to be a fundamental prerequisite for the rapid advancement of large-scale machine learning models.
@@
 In the field of molecular modeling, existing benchmarks exhibit two significant limitations.
-Firstly, they are intrinsically domain-specific, concentrating on isolated sub-fields rather than encompassing a variety of atomic systems.
+Firstly, they are intrinsically domain-specific, concentrating on isolated sub-fields rather than encompassing a variety of atomistic systems.
 For instance, datasets such as QM9[23] and MD17[24] are used to benchmark molecular property predictions and molecular dynamics (MD) trajectories of small molecules, respectively.
@@
 For instance, conventional evaluation metrics based on static test sets may not adequately capture the true performance of a model in tasks requiring physically meaningful energy landscapes[27].
-Specifically, non-conservative models - where atomic forces are directly inferred from neural networks rather than obtained from the gradient of the predicted energy[27] — can exhibit high apparent accuracy but struggle in applications demanding strict energy conservation, such as MD simulations[28].
+Specifically, non-conservative models – where atomic forces are directly inferred from neural networks rather than obtained from the gradient of the predicted energy[28] – can exhibit high apparent accuracy but struggle in applications demanding strict energy conservation, such as MD simulations[29].
 The MLIP-Arena benchmark[30] is a step in the direction toward bridging this gap, emphasizing the practical usability of LAMs in tasks such as MD stability and physical property predictions.
-However, it places less emphasis on evaluating a model’s capacity to generalize across diverse atomic systems or adapt to tasks beyond its training scope, both of which are essential for assessing the performance of LAMs in real scientific discovery.
+However, it places less emphasis on evaluating a model’s capacity to generalize across diverse atomistic systems or adapt to tasks beyond its training scope, both of which are essential for assessing the performance of LAMs in real scientific discovery.
 To address these limitations, we introduce LAMBench, a comprehensive benchmark system designed to rigorously evaluate LAMs across domains, simulation regimes, and application scenarios.
-Employing LAMBench, we evaluated the performance of eight prominent LAMs released prior to April 1, 2025, uncovering a significant discrepancy between these models and the universal PES.
+Utilizing LAMBench, we assessed the performance of ten leading LAMs released before August 1, 2025, excluding the UMA models[17] due to licensing restrictions[31].
+Our analysis revealed a substantial discrepancy between these models and the universal PES.
 Our findings suggest that enhancing LAM performance requires simultaneous training with data from a diverse array of research domains.
@@
 Figure 1: The schematic plot of the LAMBench benchmark.
-The LAMBench system is designed to benchmark diverse Large Atomic Models (LAMs) across multiple tasks within a high-throughput workflow, with automation integral to task calculation, result aggregation, analysis, and visualization, as depicted in Figure 1.
-The implementation details of LAMBench are elaborated in Section IV.2.
+The LAMBench system is designed to benchmark diverse Large Atomistic Models (LAMs) across multiple tasks within a high-throughput workflow, with automation integral to task calculation, result aggregation, analysis, and visualization, as depicted in Figure 1.
+The implementation details of LAMBench are elaborated in Section IV.3.
 Key to the system’s effectiveness are the design of the benchmark tasks and the methodologies employed for result interpretation.
 These benchmark tasks are developed to assess three fundamental capabilities of an LAM: generalizability, adaptability, and applicability.
-Generalizability pertains to the accuracy of an LAM when utilized as a universal potential across a diverse range of atomic systems.
+Generalizability pertains to the accuracy of an LAM when utilized as a universal potential across a diverse range of atomistic systems.
 Adaptability denotes the LAM’s capacity to be fine-tuned for tasks beyond potential energy prediction, with particular emphasis on structure-property relationship tasks.
 Applicability, on the other hand, concerns the stability and efficiency of deploying LAMs in real-world simulations.
+In this study, all tasks are defined, and all results are obtained using LAMBench version v0.3.1.
 Generalizability refers to the accuracy of an LAM on datasets that are not included in the training set.
@@
 Some researchers define OOD data as those exploring different configurational spaces[32, 33], while others highlight differences in chemical space as critical[26, 34].
-In this study, we adopt a practical approach by considering OOD test datasets as downstream datasets designed to address specific scientific challenges, such as training a machine learning potential model for simulating carbon deposition on metal surfaces[33].
+In this study, we adopt a practical approach by considering OOD test datasets as downstream datasets designed to address specific scientific challenges[35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46].
 These scenarios are most compatible with the downstream applications of LAMs.
 In LAMBench, the generalizability is quantitatively assessed using two types of tasks, the force field task and the property calculation task.
-The force field task assesses the accuracy of LAMs in predicting energy, force, and the virial tensor (when periodic boundary conditions are applied) using 17 datasets that have been utilized in the literature to train machine learning potentials for addressing scientific challenges.
+The force field task measures the accuracy of LAMs in predicting energy, force, and the virial tensor (under periodic boundary conditions) using twelve downstream datasets employed in the literature to train machine learning potentials for solving scientific problem, without any fine-tuning.
 Detailed information on these datasets is provided in LABEL:tab:datasets.
-These test datasets encompass five distinct scientific domains: Inorganic Materials, Catalysis, Reactions, Small Molecules, and Biomolecules and Supramolecules.
-Given that energy labels calculated via DFT can vary by an arbitrary constant due to variations in pseudopotential selection and software implementations, LAMs are consistently used to predict the energy difference between the label and a dummy model that estimates potential energy solely based on the chemical formula.
+These test datasets encompass three distinct scientific domains: Inorganic Materials, Catalysis, and Molecules.
+It is important to note that the majority of LAMs are trained on datasets labeled with PBE/PBE+U exchange-correlation (XC) functionals, which may not align with the XC functional used in the test datasets.
+To address this mismatch when evaluating the generalizability of LAMs, we relabeled five datasets at the PBE level, as detailed in the Methods section Section IV.
+Dispersion corrections were added through post-prediction process with the DFTD3-ASE interface[47, 48], when necessary.
+Minor mismatches in pseudopotentials, basis sets, and software implementations within DFT, which are less significant than the XC functional discrepancies, were not addressed in the current benchmark.
+Given that energy labels calculated via DFT can vary by an arbitrary constant, LAMs are consistently used to predict the energy difference between the label and a dummy model that estimates potential energy solely based on the chemical formula.
 In contrast, force and virial predictions are directly obtained from the LAMs, as the dummy model invariably yields null predictions for force and virial.
-To evaluate the generalizability of LAMs, it is crucial to establish a comprehensive performance indicator that integrates all the energy, force, and virial errors calculated for each configuration across every dataset and domain.
+Establishing a comprehensive performance indicator that synthesizes energy, force, and virial errors for each configuration across all datasets and domains presents a complex challenge.
 An arithmetic mean of the error data is inappropriate due to the differing units of the quantities involved.
-To address this challenge, we introduce a dimensionless error metric, $\bar{M}^{m}_{\mathrm{FF}}$, with detailed information available in Section IV.3.
-The error metric is structured so that a dummy model yields a value of $\bar{M}^{m}_{\mathrm{FF}}=1$, whereas an ideal model that perfectly aligns with DFT labels achieves a metric of $\bar{M}^{m}_{\mathrm{FF}}=0$.
+To address this issue, we introduce a dimensionless error metric, $\bar{M}^{m}_{\mathrm{FF}}$, which essentially averages the relative errors in predictions across all datasets.
+Detailed information on this metric can be found in Section IV.4.
+The metric is structured so that a dummy model yields a value of $\bar{M}^{m}_{\mathrm{FF}}=1$, whereas an ideal model that perfectly aligns with DFT labels achieves a metric of $\bar{M}^{m}_{\mathrm{FF}}=0$.
 A lower value of this error metric signifies enhanced generalizability.
+In the error metric, the relative error is preferred over the absolute error to better reflect a model’s accuracy in relation to the task’s difficulty.
+For example, the standard deviations in energy labels for the Lopanitsyna2023Modeling and MD22 datasets are 511 meV/atom and 9 meV/atom, respectively.
+This considerable difference in scales indicates the greater diversity of configurations in the Lopanitsyna2023Modeling dataset.
+If a model exhibits the same energy RMSE of 9 meV/atom for both datasets, it achieves a 2 % error in the Lopanitsyna2023Modeling dataset, while demonstrating no accuracy in the MD22 dataset.
+Therefore, given the substantial differences between datasets, the relative error offers a more meaningful measure of model performance.
 Achieving high accuracy in the force field task does not necessarily ensure enhanced performance in property calculations when utilizing the LAM as the force field[27].
@@
 Consequently, models lacking smoothness up to the second-order derivatives may not consistently yield accurate predictions, even if they perform well in force field tasks.
-Given the significant variation in investigated properties across different research domains, LAMBench offers a flexible and modular design (see Section IV.2).
+Given the significant variation in investigated properties across different research domains, LAMBench offers a flexible and modular design (see Section IV.3).
 This design facilitates the implementation of property calculation benchmark tasks, enabling them to be tailored to the specific requirements of diverse research fields.
-In this study, we have adapted the MDR Phonon benchmark[34] to evaluate the performance of LAM models in computing key phonon properties, including maximum phonon frequency ($\omega_{\max}$), entropy ($S$), free energy ($F$), and heat capacity at constant volume ($C_{V}$).
-Additionally, we have utilized the TorsionNet500 benchmark to assess the capability of LAMs in calculating the torsion energy profile of typical drug-like fragments[35].
-The Phonon and TorsionNet500 benchmarks are classified as properties of interest within the Inorganic Materials and Small Molecules domains, respectively.
-These adaptations serve as prototypes for evaluating downstream generalizability on domain-specific property calculations.
-Further tasks targeting Reactions, Catalysis and other research domains are under active development to align with the evolving capabilities of LAMs.
-Similar to the approach taken in the force field task, we have introduced a dimensionless error metric, denoted as $\bar{M}^{m}_{\mathrm{PC}}$.
-This metric is designed to assess the generalizability of an LAM in the property calculation task.
-For comprehensive details, please refer to Section IV.3.
+In this study, we include property calculation tasks across the aforementioned research domains.
+For the Inorganic Materials domain, we adapt the MDR Phonon benchmark [49] to assess the performance of LAM models in computing essential phonon properties, including maximum phonon frequency ($\omega_{\max}$), entropy ($S$), free energy ($F$), and heat capacity at constant volume ($C_{V}$).
+Additionally, we employ the elasticity benchmark [50, 51] to evaluate the accuracy of LAM models in calculating the shear modulus (GVHR) and bulk modulus (KVHR).
+For the Molecules domain, we use the TorsionNet500 benchmark[52] to evaluate LAM performance in calculating torsion energy profiles of typical drug-like fragments, and the Wiggle150 benchmark[53] to assess energy profiles of highly strained conformers.
+For the Catalysis domain, we employ the OC20NEB-OOD dataset[54] to benchmark the ability of LAMs to predict energy barriers for three reaction types—transfer, dissociation, and desorption—using DFT-optimized structures along the NEB trajectories.
+For benchmarks in the Inorganic Materials domain, the PBE/PBE+U XC functional was adopted to compute the reference predictions.
+For the Molecules and Catalysis domains, reference data are obtained at the CCSD(T) and RPBE levels, respectively, which are considered to be more accurate than the PBE XC functional within their respective domains.
+The incorporation of these property calculation tasks serves as a prototype for assessing downstream generalizability in domain-specific property calculations.
+Additional tasks can be seamlessly integrated to align with the evolving capabilities of LAMs.
+In line with the approach used in the force field task, we introduce a dimensionless error metric, denoted as $\bar{M}^{m}_{\mathrm{PC}}$, to evaluate the generalizability of an LAM in property calculation tasks.
+For comprehensive details, please refer to Section IV.4.
 Adaptability assesses the ability with which a pre-trained LAM can be fine-tuned for tasks beyond its initial training scope, with a particular focus on establishing structure-property relationships.
@@
 LAMs are designed to be applicable to tasks involving systems of varying sizes, from evaluating the conformational energy of molecules composed of dozens of atoms to large-scale simulations of viruses involving millions of atoms[57].
-The efficiency of LAMs—measured as the computational time consumed per atom while determining the energy, forces, and virial of atomic systems—is significantly influenced by the system size and the extent to which the many-core parallelism of modern GPUs is utilized.
+The efficiency of LAMs—measured as the computational time consumed per atom while determining the energy, forces, and virial of atomistic systems—is significantly influenced by the system size and the extent to which the many-core parallelism of modern GPUs is utilized.
 For relatively small systems, atom-wise efficiency is often reduced due to limited opportunities for parallelization, even though they require less overall computational time for evaluation.
@@
 However, such systems are not ideal benchmarks for assessing the efficiency of LAMs, as they are generally computationally intensive, and multi-GPU parallelism is not supported by all the LAMs.
-To address the significant size-dependency issues when measuring efficiency in small systems, we concentrate on atomic systems that are large enough to achieve converged efficiency yet small enough to be computed without the need for multi-GPU parallelism.
-To this end, we randomly sample 1,600 structures from the Inorganic Materials and Catalysis domains in the aforementioned force field task test sets.
-Each structure is subjected to periodic boundary conditions and duplicated to contain between 800 and 1,000 atoms.
-This range is sufficiently large to approach the convergence regime while remaining within the memory constraints of most models, thereby avoiding out-of-memory (OOM) errors.
-Other domains, such as Small Molecules, are excluded from this evaluation, as non-periodic systems generally do not pose sufficient computational demands to effectively stress the inference capabilities of LAMs.
+To address the significant size-dependency issues when measuring efficiency in small systems, we concentrate on atomistic systems that are large enough to achieve converged efficiency yet small enough to be computed without the need for multi-GPU parallelism.
+To this end, we randomly sample 1,000 structures from the Inorganic Materials and Catalysis domains in the aforementioned force field task test sets.
+Each structure is subjected to periodic boundary conditions and is replicated by a multiplier, dynamically determined using a binary search algorithm.
+This approach, with an upper limit of 1,000 atoms for efficiency, ensures that measurements are conducted in or near the convergence regime.
+The Molecules domain is excluded from this evaluation, as its systems contain only a few dozen atoms and therefore do not impose sufficient computational demands to meaningfully stress the inference capabilities of LAMs.
 To facilitate the comparison of efficiency across models, we propose a dimensionless efficiency metric, denoted as $M_{\mathrm{E}}^{m}$.
-This metric is defined as the normalized inverse of the averaged inference time measured for the 1,600 structures.
-The detailed definition is provided in Section IV.4.
+This metric is defined as the normalized inverse of the average inference time measured for 900 out of the 1,000 randomly sampled structures.
+A detailed definition is provided in Section IV.5.
 A value of $M_{\mathrm{E}}^{m}=1$ corresponds to an efficiency equivalent to a reference value of 0.01 $(\mathrm{\mu s/atom})^{-1}$, with higher value indicating greater efficiency.
-It is crucial to note that the calculator is fully warmed up by processing 400 structures prior to recording the inference time for the 1,600 structures.
-We assessed the stability of LAMs by examining energy conservation in NVE ensemble (microcanonical ensemble) MD simulations across nine atomic systems.
-These systems were randomly selected from diverse domains, including four periodic structures from the Materials Project[39], three molecular systems from the SPICE2 dataset[40], and two catalytic surface systems from the OC2M dataset[41], as detailed in Supporting Information Figure S-1.
+It is important to note that the calculator is fully warmed up by processing 100 structures, which are excluded from the 1,000 randomly sampled structures, before recording the inference time on the remaining 900 structures.
+We assessed the stability of LAMs by examining energy conservation in NVE ensemble (microcanonical ensemble) MD simulations across nine atomistic systems.
+The extent of drift observed over time serves as a practical indicator of a model’s stability, specifically in maintaining a finite system energy in extended MD trajectories.
+These test systems were randomly selected from diverse domains, including four periodic structures from the Materials Project[58], three molecular systems from the SPICE2 dataset[59], and two catalytic surface systems from the OC2M dataset[60], as detailed in Supporting Information Figure S-1.
 For each system, a 10 ps NVE-MD simulation was conducted with a timestep of 1 fs.
@@
 Stability is assessed via the instability metric, $M^{m}_{\mathrm{IS}}$, which is calculated based on total energy drift, where smaller values denote improved stability.
-A detailed definition is provided in Section IV.4.
+A detailed definition is provided in Section IV.5.
 Table 1: Summary of datasets utilized in force field generalizability assessments.
 The table details the domain of each test dataset, available labels (E for energy, F for force, and V for virial), number of configurations (frames), exchange-correlation (XC) functional applied for labeling, and additional descriptions.
-Details on data cleaning procedures can be found in Supporting Information Section S-3.
-Name Domain Labels Frames Level of XC Description
-Lopanitsyna2023Modeling_A[43] Inorganic Materials EFV 450 PBE-sol Commonly known as HEA25S. A dataset of high-entropy alloy surfaces, focusing on d-block transition metals covering 25 elements.
-Lopanitsyna2023Modeling_B[44] Inorganic Materials EFV 25628 PBE-sol Commonly known as HEA25.
-A dataset of high-entropy alloy bulk structures, focusing on d-block transition metals covering 25 elements.
-Dai2024Deep[45] Inorganic Materials EFV 2842 PBE A dataset of high-entropy transition metal diboride (HEMB2) and carbide (HEMC) ceramics.
-Torres2019Analysis[46] Inorganic Materials EF 4378 PBE A dataset of Ca-bearing minerals, focusing on silicates and carbonates.
-Sours2023Applications[47] Inorganic Materials EF 17414 PBE-D3(BJ) Dataset consisting of 219 different pure silica zeolite topologies.
-Batzner2022equivariant[48] Inorganic Materials EF 7499 PBE Down-sampled dataset containing lithium phosphate amorphous glass used by NequIP graph neural network models as evaluation configurations.
-WBM_25k[49] Inorganic Materials E 25696 PBE A subset of the WBM dataset randomly down-sampled to 25,696 frames.
-Subalex_9k[50] Inorganic Materials EFV 9157 PBE A subset of sAlexandria dataset, randomly down-sampled to 9,157 frames.
-ANI-1x[51] Small Molecules EF 8861 $\omega$B97X / 6-31G* A down-sampled dataset from the training data of the ANI-1x potential, containing organic molecule structures.
-Zhang2024Active[33] Catalysis EF 452 PBE-D3 Dynamic simulations of carbon deposition on metal surfaces.
-Zhang2019Bridging[52] Catalysis EF 9842 PBE Interaction of carbon dioxide with a movable Ni(100) surface.
-Vandermause2022Active[53] Catalysis EFV 250 PBE Direct simulation of hydrogen turnover on Pt(111) catalyst surfaces.
-Villanueva2024Water[54] Catalysis EF 14859 PBE-D3 Selective CO2 hydrogenation to methanol over oxide catalysts.
-Guan2022Benchmark[55] Reactions EF 17412 $\omega$B97X-V / cc-pVTZ Dataset of hydrogen combustion reactions covering 19 reaction channels.
-Gasteiger2020Fast[56] Reactions EF 9480 revPBE-D3 / def2-TZVP Validation set from molecular collision experiments with small organic molecules.
-MD22[57] Biomolecules EF 2122 PBE+MBD / tighta Molecular dynamics (MD) trajectories of 42-atom tetrapeptide Ac-Ala3-NHMe.
-AIMD-Chig[58] Biomolecules EF 19800 M06-2X / 6–31G* A down-sampled MD dataset containing conformations of chignolin protein.
-a The “tight” setting corresponds to the default high-accuracy mode in FHI-aims.[42]
+Name Domain Labels Frames Description
+Mazitov2024Surface[35] Inorganic Materials EFV 422 A down-sampled dataset of HEA25S, for high-entropy alloy surfaces, focusing on d-block transition metals covering 25 elements.
+Lopanitsyna2023Modeling[36] Inorganic Materials EFV 900 A down-sampled dataset of HEA25, for high-entropy alloy bulk structures, focusing on d-block transition metals covering 25 elements.
+Torres2019Analysis[37] Inorganic Materials EF 1000 A down-sampled dataset of Ca-bearing minerals, focusing on silicates and carbonates.
+Sours2023Predicting[38] Inorganic Materials EF 1000 A down-sampled Dataset consisting of 219 different pure silica zeolite topologies.
+Batzner2022equivariant[39] Inorganic Materials EF 1000 A down-sampled dataset containing lithium phosphate amorphous glass used by NequIP graph neural network models as evaluation configurations.
+Gao2025Spontaneous[40] Inorganic Materials EFV 1000 A down-sampled Bilayer 2D MoS2 structures.
+ANI-1x[41] Molecules EF 997 A down-sampled dataset from the training data of the ANI-1x potential, containing organic molecule structures.
+MD22[42] Molecules EF 1000 A down-sampled dataset from MD22, where molecular dynamics (MD) trajectories of 42-atom tetrapeptide Ac-Ala3-NHMe, docosahexaenoic acid, Stachyose, DNA base pair (AT-AT), DNA base pair (AT-AT-CG-CG).
+AIMD-Chig[43] Molecules EF 942 A down-sampled MD dataset containing conformations of chignolin protein.
+Zhang2019Bridging[44] Catalysis EF 1000 A down-sampled dataset for interaction of carbon dioxide with a movable Ni(100) surface.
+Vandermause2022Active[45] Catalysis EFV 250 A down-sampled dataset for direct simulation of hydrogen turnover on Pt(111) catalyst surfaces.
+Villanueva2024Water[46] Catalysis EF 1000 A down-sampled dataset for selective CO2 hydrogenation to methanol over oxide catalysts.
 II.2 Benchmark results
@@
 This leaderboard aims to strike a balanced comparison between key performance aspects and is intended to offer insights into model suitability for applications relevant to real-world scientific discovery.
+To reduce discrepancies in the comparisons between multi-task and single-task trained models, we employed the MPtrj task head for the DPA-3.1-3M and DPA-2.4-7M models, and the MPA modality for the SevenNet-MF-ompa model, unless stated otherwise.
 Table 2: The LAMBench Leaderboard.
@@
 $\bar{M}^{m}_{\mathrm{FF}}\downarrow$ $\bar{M}^{m}_{\mathrm{PC}}\downarrow$ $M_{\mathrm{E}}^{m}\uparrow$ $M^{m}_{\mathrm{IS}}\downarrow$
-DPA-2.4-7M 0.265 0.208 0.614 0.039
-GRACE-2L-OAM 0.340 0.262 0.678 0.309
-SevenNet-l3i5 0.355 0.240 0.279 0.036
-MACE-MPA-0 0.356 0.291 0.291 0.000
-Orb-v2 0.356 0.560 1.343 2.649
-SevenNet-MF-ompa 0.358 0.300 0.088 0.000
-MatterSim-v1-5M 0.389 0.280 0.388 0.000
-MACE-MP-0 0.405 0.341 0.291 0.089
+DPA-3.1-3M 0.175 0.322 0.261 0.572
+Orb-v3 0.215 0.414 0.396 0.000
+DPA-2.4-7M 0.241 0.342 0.617 0.039
+GRACE-2L-OAM 0.251 0.404 0.639 0.309
+Orb-v2 0.253 0.601 1.341 2.649
+SevenNet-MF-ompa 0.255 0.455 0.084 0.000
+MatterSim-v1-5M 0.283 0.467 0.393 0.000
+MACE-MPA-0 0.308 0.425 0.293 0.000
+SevenNet-l3i5 0.326 0.397 0.272 0.036
+MACE-MP-0 0.351 0.472 0.296 0.089
 Figure 2: Dimensionless error metrics for generalizability tasks across different domains.
-(a) Dimensionless error metrics for force field tasks across five distinct domains.
-(b) Property-calculation tasks: the Small Molecules domain is assessed using the TorsionNet500 benchmark, while the Inorganic Materials domain is evaluated using the phonon MDR benchmark.
-To reassess DPA-2.4-7M with enhanced XC functional alignment, referred to as DPA-2.4-7M-bestXC, we selected task heads trained on Yang2023ab for the ANI-1x and AIMD-Chig datasets.
-For the TorsionNet500 benchmark, we adopted the task head trained on the SPICE2 dataset.
-For the Sours2023Applications dataset, we applied a D3(BJ) dispersion correction.
+(a) Dimensionless error metrics for force field tasks across three distinct domains.
+(b) Dimensionless error metrics for property-calculation tasks.
+To reassess DPA-3.1-3M with enhanced domain alignment, referred to as DPA-3.1-3M-DomainMatch, we employed task heads trained on SPICE2 for the Molecules domain, OMAT24 for the Inorganic Materials domain, and OC20M for the Catalysis domain.
 In the force field generalizability task, as illustrated in Table 2, performance is evaluated using the dimensionless error $\bar{M}_{\mathrm{FF}}^{m}$.
-Notably, DPA-2.4-7M demonstrates substantially greater generalizability compared to other LAMs.
-Interestingly, aside from MatterSim-v1-5M and MACE-MP-0, most models exhibit only minor differences in generalizability.
+Notably, DPA-3.1-3M demonstrates substantially greater generalizability compared to other LAMs.
 To further investigate the generalizability of these LAMs across specific domains, we present the domain-wise dimensionless error metric in Figure 2(a).
-In general, most LAMs exhibit relatively low inference errors in the domains of Inorganic Materials, Reactions, and Small Molecules, while showing inferior performance in the domains of Catalysis and Biomolecules and Supramolecules.
-In the Reactions domain, DPA-2.4-7M, Orb-v2, and SevenNet-MF-ompa achieve the lowest errors, with MatterSim-v1-5M following closely behind.
-For Small Molecules, DPA-2.4-7M and Orb-v2 rank first and second, respectively, outperforming all other models in this domain.
-In the Inorganic Materials domain, the comparative performance of the models shows significant alignment with their rankings in the Matbench Discovery benchmark, accessed on April 1st, 2025.
-The models are ranked by their highest F1 scores as follows, excluding DPA-2.4-7M: SevenNet-MF-ompa $>$ GRACE-2L-OAM $>$ Orb-v2 $>$ MatterSim-v1-5M $>$ MACE-MPA-0 $>$ SevenNet-l3i5 $>$ MACE-MP-0.
+In general, most LAMs exhibit relatively low inference errors in the domains of Molecules and Inorganic Materials, while showing inferior performance in the domain of Catalysis.
+For Molecules, DPA-3.1-3M ranks first, closely followed by DPA-2.4-7M and Orb-v3, and outperforms all other models in this domain.
+In the Inorganic Materials domain, the comparative performance of the models shows a similar pattern to their rankings in the Matbench Discovery benchmark, accessed on August 1st, 2025.
+These models, excluding DPA2-2.4-7M, are ranked by their lowest MAE in the following order: SevenNet-MF-ompa << GRACE-2L-OAM << DPA-3.1-3M << Orb-v3 << MatterSim-v1-5M << MACE-MPA-0 << Orb-v2 << SevenNet-l3i5 << MACE-MP-0.
 This coherence further substantiates the reliability and robustness of both the LAMBench and Matbench Discovery benchmark frameworks[25] in evaluating LAMs for Inorganic Materials force field tasks.
-Within the Catalysis domain, DPA-2.4-7M again achieves the lowest error, significantly surpassing the other models.
-In the Biomolecules and Supramolecules domain, the DPA-2.4-7M model achieves the highest generalizability, distinctly outperforming all other LAMs.
-The remaining models show comparable performances, except for MACE-MP-0 and MACE-MPA-0, which fall noticeably behind.
-Table 2 reveals that MACE-MPA-0, characterized by a larger parameter set and trained on an extensive dataset, exhibits a dimensionless error 0.05 lower than MACE-MP-0, suggesting enhanced generalizability.
-This trend is corroborated by the domain-wise error metrics in Figure 2 (a), which demonstrate superior performance across most domains, with the exception of Catalysis, where a minor decline is noted.
-The Inorganic Materials domain shows the most significant improvement, with an approximately 30 percent reduction in error.
-Meanwhile, SevenNet-MF-ompa delivers overall performance on par with its lighter counterpart, SevenNet-l3i5; however, its enhanced accuracy in the Inorganic Materials and Reactions domains comes at the cost of reduced generalizability elsewhere, indicating potential overfitting to the Inorganic Materials domain.
-The DPA-2.4-7M model demonstrated the best overall performance across all investigated domains.
-This success can be attributed to its training on 31 datasets, detailed in Supporting Information LABEL:tab:si-datasets, which span multiple domains using a multitask training strategy[16, 4].
-In this approach, a common feature extraction structure, the descriptor, is linked to multiple task heads.
-Each task head is trained on one of the 31 datasets, while the common descriptor is simultaneously trained on all datasets.
-Consequently, the descriptor is designed to extract the shared knowledge from all the training datasets.
-For a fair comparison, the task head trained with the MPtrj dataset, which serves as a subset of the training data for all LAMs, is consistently used to assess the generalization error.
-Thus, the DPA-2.4-7M model’s exceptional performance in the Small Molecules, Biomolecules and Supramolecules, and Catalysis domains can be attributed to the descriptor, which was trained using diverse datasets from domains such as Small Molecules (including the SPICE2 and Yang2023ab datasets) and Catalysis (including the OC20 and OC22 datasets).
-It is important to note that certain test datasets are labeled with an exchange-correlation (XC) functional that differs from the one used for labeling the training datasets, such as MPtrj, for the LAMs.
-For instance, in the domain of Small Molecules, the ANI test set is labeled by $\omega$B97X/6-31G* functional.
-To further investigate the impact of XC functional mismatch between the model’s training data and the reference functional used in generalizability tests, we reassessed the generalization error of the DPA-2.4-7M model using task heads that align as closely as possible with the XC functional employed for labeling the test datasets.
-This alignment resulted in a substantial reduction in generalization error across several domains.
-Specifically, the domain error metric decreased from 0.19 to 0.09 for Small Molecules and from 0.25 to 0.17 for Biomolecules and Supramolecules (see Figure 2 (a)).
-These findings suggest that XC functional mismatch significantly contributes to the generalization error of the LAM model, highlighting the necessity of training models with multiple XC functional fidelities[59].
+Within the Catalysis domain, DPA-3.1-3M again achieves the lowest error, significantly surpassing the other models.
+Table 2 indicates that MACE-MPA-0, which features an expanded set of parameters and is trained on a more comprehensive dataset, achieves a dimensionless error of 0.308.
+It is 0.043 lower than the error of MACE-MP-0, which stands at 0.351, suggesting improved generalizability.
+This finding is further supported by the domain-specific error metrics presented in Figure 2(a), where MACE-MPA-0 demonstrates superior performance across all three domains.
+Notably, in the Inorganic Materials domain, there is an approximate 25% reduction in error.
+Similarly, the SevenNet-MF-ompa model shows a comparable improvement over its lighter version, SevenNet-l3i5.
+The DPA-3.1-3M model[61] demonstrates the best overall performance across all investigated domains.
+This achievement is likely due to its multi-task training strategy[16, 4, 17] that train the model on 31 datasets spanning multiple domains, detailed in Supporting Information LABEL:tab:si-datasets.
+The model’s outstanding performance in the Molecules and Catalysis domains can be attributed to the descriptor, which was trained using diverse datasets from domains such as Molecules (including the SPICE2 and Yang2023ab datasets) and Catalysis (including the OC20 and OC22 datasets).
+The DPA-2.4-7M model, trained using the same multitask strategy, ranks third, showing a noticeable gap from DPA-3.1-3M and Orb-v3, largely due to its limited capacity.
 The generalizability of LAMs in property calculation tasks is assessed and summarized in Table 2.
 Compared to force field tasks, property calculation tasks offer greater discriminative power among models.
-The DPA-2.4-7M model again achieves the best performance, with the lowest error of 0.208—representing nearly a two-thirds reduction in error relative to the worst-performing model.
-The domain-wise error metric is shown in Figure 2(b), with the detailed raw measurements available in Supporting Information Tables S-3 and S-4.
-In the Inorganic Materials domain, where the MDR phonon benchmark is conducted, the SevenNet-MF-ompa, MatterSim-v1-5M, and GRACE-2L-OAM models demonstrate superior performance, with MACE-MPA-0 following closely.
-Notably, the conservative models exhibit significantly lower errors compared to the non-conservative Orb-v2 model, which directly predicts force.
-Accurate phonon predictions necessitate the calculation of the force constant matrix, representing the second-order derivative of a force field.
-Therefore, conservativeness and smoothness are crucial for achieving precise phonon predictions, consistent with previous findings in Ref.[26].
-Additionally, the performance difference between MACE-MP-0 and its more expressive counterpart, MACE-MPA-0, suggests that enhanced generalizability in force field prediction tasks can lead to improved performance in a property prediction task.
-In the Small Molecules domain, where the TorsionNet500 benchmark is conducted, the dimensionless error values are notably higher compared to those in the Inorganic Materials domain.
-The multi-task trained LAM, DPA-2.4-7M, demonstrates the lowest error at 0.35, outperforming other models primarily trained on datasets within the Inorganic Materials domain.
-Utilizing the task head trained with the SPICE2 dataset, which employs a closely aligned XC as the reference DFT method, reduces the error to 0.19.
-Despite these improvements, performance remains significantly lower than that of domain-specific models such as MACE-OFF23[60], which achieves a dimensionless error of 0.034 on TorsionNet500 benchmark.
-The discrepancy between the performance of LAMs and MACE-OFF23 highlights the necessity of focusing on domains beyond Inorganic Materials to enhance generalizability in property prediction tasks.
-This also indicates that current LAMs remain considerably less accurate than domain-specific models and are still distant from achieving the capabilities of a universal PES model.
-The adaptation of the pretrained DPA-2.4-7M model across eight Matbench property regression tasks is detailed in Table 3.
+The DPA-3.1-3M model again achieves the best performance, with the lowest error of 0.322—representing nearly a fifty percent reduction in error relative to the worst-performing model.
+The domain-wise error metrics are presented in Figure 2(b), with the detailed raw measurements available in Supporting Information Tables S-3, S-4, S-5, S-6 and S-7.
+In the Inorganic Materials domain, where the MDR phonon and elasticity benchmarks are conducted, the conservative models exhibit significantly lower errors compared to the non-conservative Orb-v2 model, which directly predicts force.
+Accurate phonon and elasticity predictions necessitate the calculation of the force constant matrix, representing the second-order derivative of a force field.
+Therefore, conservativeness and smoothness are crucial for achieving precise predictions, consistent with previous findings in Ref.[27].
+Additionally, the performance difference between MACE-MP-0 (0.14) and its more expressive counterpart, MACE-MPA-0 (0.06), suggests that enhanced generalizability in force field prediction tasks can lead to improved performance in a property prediction task.
+In both benchmarks, the Omat24 task head of DPA-3.1-3M, denoted as DPA-3.1-3M-DomainMatch, demonstrates improved performance over the MPtrj task head, underscoring the benefits of more extensive training data.
+[…]
+In the Catalysis domain, where the OC20NEB-OOD benchmark is conducted, all LAMs exhibit limited generalizability.
+The top-performing model, DPA-3.1-3M-DomainMatch, utilizing the OC20M task head, still presents a relatively high dimensionless error of 0.53.
+In contrast, domain-specific models such as EquiformerV2-31M-S2EF-OC20-All+MD[64, 65] and eSCN-L6-M3-Lay20-S2EF-OC20-All+MD[66, 65] achieve substantially lower dimensionless errors of 0.31 and 0.33, respectively.
+As illustrated in Table S-7, LAMs achieve notably lower errors when predicting the reaction energy ($dE$) as opposed to the reaction barrier ($E_{a}$).
+This discrepancy indicates that improving the performance of LAMs necessitates increased focus on the transition states.
+The adaptation of the pretrained DPA-3.1-3M and DPA-2.4-7M models across eight Matbench property regression tasks is detailed in Table 3.
 The fine-tuned model consistently surpasses the model with the same architecture trained from scratch in terms of accuracy.
-Despite the improvements achieved by the finetuned DPA-2.4-7M model, a discrepancy remains between its performance and the results showcased on the Matbench leaderboard.
 Notably, a property predictor adapted from the pretrained MatterSim[67] achieves performance comparable to leading task-specific models on the Matbench leaderboard.
 Furthermore, JMP, pretrained on the OC20, OC22, ANI-1x, and Transition-1x datasets[16], significantly surpasses existing models, achieving state-of-the-art accuracy.
+Although a performance gap remains between DPA-3.1-3M and the state-of-the-art (SOTA) results, its superior performance compared to DPA-2.4-7M demonstrates that better generalizability in force field prediction tasks also translates to improved adaptability in property fine-tuning tasks.
 This underscores the critical importance of pretraining for achieving superior model accuracy in downstream property prediction tasks and reinforces the broader vision of LAMs as versatile, high-performing surrogates capable of addressing a wide range of scientific challenges.
-Table 3: Adaptability Test: The accuracy of property fine-tuning of DPA-2.4-7M across eight Matbench regression tasks.
-Task (Unit) DPA-2.4-7M From scratch DPA-2.4-7M Finetune11footnotemark: 1 Matbench[37] leaderboard MatterSim[61] JMP[16]
-MP $E_{\text{form}}$ (meV/atom) 30.1 24.1 17.0 - 10.1
-MP Gap (eV) 0.256 0.206 0.156 0.129 0.091
-JDFT2D (meV/atom) 43.68 34.09 33.19 32.76 29.94
-Phonons (cm-1) 43.55 34.00 28.76 26.02 20.57
-Dielectric (unitless) 0.418 0.300 0.271 0.252 0.249
-Log KVRH ($\log_{10}\text{GPA}$) 0.061 0.052 0.049 0.049 0.045
-Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.068 0.067 0.061 0.059
-Perovskites (eV/unitcell) 0.054 0.032 0.027 - 0.026
+Table 3: Adaptability Test: The accuracy of property fine-tuning of DPA models across eight Matbench regression tasks. 11footnotemark: 1
+Task (Unit) DPA-2.4-7M From scratch DPA-2.4-7M Finetune DPA-3.1-3M From scratch DPA-3.1-3M Finetune Matbench[56] leaderboard MatterSim[67] JMP[16]
+MP $E_{\text{form}}$ (meV/atom) 31.1 24.7 24.2 13.9 17.0 - 10.1
+MP Gap (eV) 0.285 0.261 0.255 0.147 0.156 0.129 0.091
+JDFT2D (meV/atom) 46.93 32.23 41.80 31.29 33.19 32.76 29.94
+Phonons (cm-1) 41.10 29.05 41.41 25.13 28.76 26.02 20.57
+Dielectric (unitless) 0.431 0.292 0.438 0.297 0.271 0.252 0.249
+Log KVRH ($\log_{10}\text{GPA}$) 0.062 0.052 0.060 0.051 0.049 0.049 0.045
+Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.066 0.079 0.061 0.067 0.061 0.059
+Perovskites (eV/unitcell) 0.061 0.047 0.043 0.035 0.027 - 0.026
 11footnotemark: 1
-Results were obtained using NVIDIA A800-SXM4-40GB GPUs with the maximum permissible batch size.
-Performance may degrade on NVIDIA L4-24GB GPUs with smaller batch sizes as illustrated in Table S-5, which are used by default in the workflow for cost management purposes.
+Results were obtained using 4 * NVIDIA L4-24GB GPUs with the maximum permissible batch size.
+For the MP $E_{\text{form}}$ and MP Gap tasks, the reported accuracy may not reflect the fully converged results due to insufficient training epochs.
 High efficiency and robust stability are equally critical as model accuracy, especially in the context of high-throughput simulations.
@@
 In contrast, Orb-v2 and DPA-2.4-7M utilize a fixed maximum number of neighbors through padding, rendering them insensitive to such variations.
-Figure 3: Distribution of inference time, normalized by the number of atoms, measured across 1,600 randomly selected configurations. Lower values indicate higher efficiency.
+Figure 3: Distribution of inference time, normalized by the number of atoms, measured across 900 randomly selected configurations. Lower values indicate higher efficiency.
 The stability of the LAMs is evaluated using the instability metric $M^{m}_{\mathrm{IS}}$, as presented in Table 2.
-The conservative models, such as MatterSim-v1-5M, exhibit minimal energy drift throughout the evaluation, indicating stable simulations.
+The conservative models, such as Orb-v3 , exhibit minimal energy drift throughout the evaluation, indicating stable simulations.
 In contrast, the non-conservative model, Orb-v2, shows instability metric several orders of magnitude larger, reflecting instability in the simulation.
+The relatively high instability metric observed in DPA-3.1-3M is due to a failed simulation, potentially caused by the relatively low training weight assigned to the MPtrj task head, as detailed in LABEL:tab:si-datasets.
+Notably, utilizing the Omat24 task head reduces the instability metric to zero.
 In general, models with a greater number of trainable parameters and trained on larger datasets exhibit higher stability, as reflected in reduced energy drift.
-This trend is evident in the MACE, and SevenNet model families as shown in Supporting Information Table S-6.
+This trend is evident in the MACE, and SevenNet model families as shown in Supporting Information Table S-8.
 III Discussion and Conclusion
-This study introduces a comprehensive benchmarking framework, designated as LAMBench, for evaluating large atomic models (LAMs).
+This study introduces a comprehensive benchmarking framework, designated as LAMBench, for evaluating large atomistic models (LAMs).
 It is designed to assess the extent to which LAMs can be used as versatile, out-of-the-box tools capable of advancing scientific discovery across a broad range of contexts.
-We emphasize three essential requirements for realizing this vision: generalizability to atomic systems across diverse research domains, adaptability to novel tasks, and applicability in real-world simulations with regard to stability and efficiency.
+We emphasize three essential requirements for realizing this vision: generalizability to atomistic systems across diverse research domains, adaptability to novel tasks, and applicability in real-world simulations with regard to stability and efficiency.
 We propose two test tasks to evaluate generalizability: the force field tasks, which assess LAMs’ accuracy in predicting energy, force, and virial (when applicable), and the property calculation tasks, which measure accuracy in computing properties of interest for specific applications using the LAM as a force field.
@@
 This metric facilitates the comparison of generalizability across different test tasks, despite significant variations in label magnitudes, and offers a clear indication of the model’s proximity to an ideal universal model.
-In this study, we benchmarked eight leading LAMs, released before April 1, 2025.
-DPA-2.4-7M consistently demonstrated superior performance in terms of generalizability compared to all other LAMs.
-In the force field generalizability test, DPA-2.4-7M was followed by GRACE-2L-OAM, SevenNet-l3i5, SevenNet-MF-ompa, MACE-MPA-0, and Orb-v2, all exhibiting remarkably similar accuracy levels.
-A trend of overfitting to the Inorganic Materials domain was observed in some LAMs.
-However, in the generalizability test for property calculation, more pronounced differences between the models are evident.
-In the phonon calculation test, the disparities are mainly attributed to the conservativeness of the models, with conservative models significantly outperforming the non-conservative ones.
-In the torsion profile calculation task within the Small Molecules domain, where most LAMs, except DPA-2.4-7M and SevenNet-l3i5, exhibit a dimensionless error of $\geq 0.5$.
-Although the XC-adapted DPA-2.4-7M model achieves the lowest generalization error of 0.19, it remains significantly higher than MACE-OFF23, a model specifically trained for the Small Molecules domain, which has an error of 0.034.
 The LAMBench results highlight a substantial gap between current LAMs and the ideal universal model, envisioned as an out-of-the-box simulation tool for scientific discovery.
 This gap is primarily due to limited generalizability beyond the Inorganic Materials domain, which is largely attributed to training datasets predominantly sourced from this domain.
-While DPA-2.4-7M employs a multi-task training scheme, simultaneously utilizing data from multiple domains to enhance generalization capabilities, its accuracy remains inferior to models specifically oriented towards individual domains.
+More specialized datasets are needed to improve cross-domain generalization, such as those containing transition states in the Catalysis domain.
+Although DPA-3.1-3M employs a multi-task training scheme, utilizing data from multiple domains to enhance its generalization capabilities and outperforming other single-task trained models in most situations, its accuracy remains inferior to models specifically tailored for individual domains.
 To bridge this gap, future research might focus on developing advanced model architectures for improved generalizability, more efficient training methods to learn shared knowledge across diverse domains, and a balanced distribution of training data across various research fields.
@@
 The capability to adjust model fidelity is proposed as an essential feature for LAMs, given that different research domains necessitate varying levels of accuracy in DFT calculations.
-For instance, the Inorganic Materials domain typically requires GGA functionals, whereas the Small Molecules domain often demands hybrid functionals.
-Notably, the force field error in the Small Molecules domain for DPA-2.4-7M decreases from 0.19 to 0.09 when switching from PBE/PBE+U to the $\omega$B97X-D functional.
-This underscores the importance of adaptable fidelity in achieving optimal performance across diverse research applications.
+For example, the Inorganic Materials domain generally necessitates GGA functionals, whereas the Molecules domain often requires hybrid or higher-level functionals.
+Training models across various XC functional fidelities offers considerable flexibility for downstream applications and should thus be actively promoted within the community [68].
 Regarding the efficiency of using LAMs in simulations, the non-conservative model Orb-v2 demonstrates significantly higher performance compared to the conservative models.
-Among the conservative models, the fastest model is 7.7 times quicker than the slowest, a disparity that is notably larger than the variations observed in generalizability.
+Among the conservative models, the fastest model is 7.6 times quicker than the slowest, a disparity that is notably larger than the variations observed in generalizability.
 This observation underscores the importance of considering the efficiency in the development of LAMs, particularly for speed-critical applications such as long-time molecular dynamics simulations.
@@
 This observation, along with the necessity for conservativeness in phonon calculation tasks, emphasizes the importance of integrating conservativeness into the design of LAMs.
-In this study, adaptability is exclusively benchmarked for the DPA-v2.4-7M model, as most implementations of LAMs are primarily confined to force field prediction and are not readily adaptable to property prediction tasks.
-The adaptability test for DPA-2.4-7M demonstrates that a pretrained model can offer substantial advantages in downstream property prediction tasks compared to models trained from scratch, particularly in data-scarce scenarios.
-Despite DPA-2.4-7M’s strong performance in generalizability tests, a noticeable gap remains between its accuracy and those of state-of-the-art property prediction models.
-This highlights a significant opportunity for further enhancing the adaptability of the DPA-2.4-7M model representing a crucial step towards the development of ideal universal models.
+In this study, adaptability is exclusively benchmarked for the DPA models, as most implementations of LAMs are primarily confined to force field prediction and are not readily adaptable to property prediction tasks.
+The adaptability test for DPA-3.1-3M and DPA-2.4-7M demonstrates that a pretrained model can offer substantial advantages in downstream property prediction tasks compared to models trained from scratch, particularly in data-scarce scenarios.
+The comparison between these two models further suggests that greater generalizability in force field prediction tasks translates to improved adaptability in property fine-tuning tasks.
+Additional research is needed to validate this correlation across other LAMs.
 Furthermore, it is advisable for LAM developers to enhance their models with the capability to adapt for property prediction.
 This advancement could open up new avenues for their application in scientific discovery.
-In the future, LAMBench should be enhanced to more accurately reflect the LAMs’ performance in real-world applications as an out-of-box simulation tool.
-This enhancement requires the collection of additional test datasets for the force field generalizability test and the incorporation of more property calculation workflows in the property generalizability test.
+Currently, LAMBench only includes twelve OOD datasets for force field prediction and five for property calculation across three research domains, supporting eight adaptability and two applicability tests.
+In the future, LAMBench should incorporate more comprehensive downstream test cases, encompassing both force field and property calculation scenarios, to more accurately reflect the performance of LAMs in real-world applications as an out-of-the-box simulation tool.
+This enhancement necessitates dynamically adjusting the force field generalizability test datasets to accommodate emerging trends in training dataset development, which may render some OOD generalizability test cases in-distribution.
+Consequently, it will be essential to replace these test cases with more challenging alternatives.
 It is important to note that the design of property calculation tasks requires careful consideration.
@@
 IV Method
-IV.1 Models
+IV.1 computational details
+Energy and force labels of ANI-1x, MD22 and AIMD-Chig sets are calculated using density functional theory (DFT) at the level of generalized gradient approximation (GGA) in the form of Perdew, Burke and Ernzerhof (PBE) functional[69][70] with 6-31G(d) basis sets in Gaussian 16 Rev.
+C.01[71].
+self-consistent field (SCF) convergence requires the root mean squared (RMS) change in the density matrix smaller than $1\times 10^{-8}$, meanwhile the maximum change in the density matrix smaller than $1\times 10^{-6}$.
+Integration grids are set to 99 radial shells per atom and 590 angular points per shell.
+Calculations for ANI-1x and MD22 systems adopt the neutral and closed-shell singlet states, which is consistent with the original datasets.
+Calculations for AIMD-Chig systems adopt the -2 charged and closed-shell singlet states, given all molecules in this set are chignolin dianions.
+The energies, interatomic forces and virials labels of the Lopanitsyna2023Modeling and Mazitov2024Surface data on were obtained from density functional theory (DFT) calculations performed using the Vienna Ab initio Simulation Package (VASP).
+The electronic structure calculations employed the generalized gradient approximation with the PBE functional for exchange-correlation effects, adopting parameter settings consistent with the Materials Project database, including a plane-wave cutoff energy of 520 eV.
+The electronic self-consistency cycle was converged to $5\times 10^{-5}$ eV.
+For Brillouin zone integration, we used the Methfessel-Paxton smearing method (ISMEAR = 1) with a smearing width of 0.05 eV, which is particularly suitable for metallic systems like high-entropy alloys.
+The k-point mesh was generated automatically based on a reciprocal density of 64, corresponding to a k-spacing of approximately 0.5 $^{-1}$, ensuring sufficient sampling accuracy while maintaining computational efficiency.
+Core-electron interactions were treated using projector-augmented wave pseudopotentials, and all calculations were performed without spin polarization to maintain consistency across the diverse compositional space of our high-entropy alloy systems.
+IV.2 Models
 To evaluate the thoroughness and discriminative power of the LAMBench benchmark, we test a series of LAMs, as listed in Table 4.
 Most models follow a single-task, single-fidelity training strategy, using datasets curated under consistent DFT settings—primarily from the domain of Inorganic Materials.
-Exceptions include SevenNet-MF-OMPA and DPA-2.4-7M.
+Exceptions include SevenNet-MF-OMPA and DPA models.
 SevenNet-MF-OMPA employs multi-fidelity training on the OMat24[72], MPtrj, and sAlex[72] datasets, achieving high accuracy despite heterogeneity in DFT settings, though still focused on inorganic systems.
-In contrast, DPA-2.4-7M adopts a multitask training approach using the OpenLAM dataset collection (Supporting Information LABEL:tab:si-datasets), which spans a broad range of chemical and material systems—including catalysis, small molecules, reactions, and biomolecules—under diverse DFT settings.
+Conversely, both the DPA-3.1-3M and DPA-2.4-7M models implement a multitask training strategy utilizing the OpenLAM dataset collection (refer to Supporting Information LABEL:tab:si-datasets).
+This collection encompasses a wide array of chemical and material systems, including inorganic materials, catalysis, small molecules, biomolecules, and reactions, all under diverse DFT settings.
 Detailed hyperparameter information for the models can be found in Supporting Information Table S-1.
-The multitask training of DPA-2.4-7M produces 31 prediction heads, each corresponding to a distinct training task within OpenLAM.
-Unless otherwise noted, the prediction head trained by the MPtrj dataset is used by default.
+Unless specified otherwise, the MPtrj task head is employed by default across all benchmarks.
 Table 4: Summary of LAMs benchmarked in this study.
@@
 Model # Parameters Training Set Direct Force Prediction Cutoff Radius (Å)
+DPA-3.1-3M[61] 3.27M OpenLAM No 6.0
 DPA-2.4-7M[4] 6.64M OpenLAM No 6.0
@@
 Orb-v2[28] 25.2M MPtrj, Alex Yes 10.0
+Orb-v3-conservative-inf-mpa[73] 25.5M MPtrj, Alex No 6.0
+SevenNet-l3i5[12] 1.17M MPtrj No 5.0
 SevenNet-MF-ompa[68] 25.7M OMat24, MPtrj, sAlex No 6.0
-SevenNet-l3i5[12] 1.17M MPtrj No 5.0
 MatterSim-v1-5M[67] 4.55M MattterSim No 5.0
-GRACE-2L-OAM[62] 15.3M MPtrj No 6.0
-IV.2 LAMBench Implementation
+GRACE-2L-OAM[74] 12.6M MPtrj No 6.0
+IV.3 LAMBench Implementation
 Benchmarking LAMs involves repeatedly performing computations using various combinations of models and test tasks, followed by aggregating and visualizing the benchmarking results.
@@
 This process enables researchers to intuitively analyze and interpret the benchmarking outcomes.
-IV.3 Generalizability metrics
+IV.4 Generalizability metrics
 In assessing the generalizability of models, the primary metrics employed are the mean absolute error (MAE) and root mean square error (RMSE) for specific predictions across test sets within various domains.
@@
 Thus, a more integrated approach to evaluation is required to effectively compare the generalizability of different models, considering the multitude of error metrics involved.
-[…]
-$\hat{M}^{m}_{k,p,i}=\frac{M^{m}_{k,p,i}}{M^{\mathrm{dummy}}_{k,p,i}}$ (1)
+In this study, we denote the error metric as $M^{m}_{k,p,i}$, where $m$ indicates the model, $k$ denotes the domain index, $p$ signifies the prediction index, and $i$ represents the test set index.
+For instance, in force field tasks, the domains include Molecules, Inorganic Materials, and Catalysis, such that $k\in\{\text{Molecules, Inorganic Materials, Catalysis}\}$.
+The prediction types are categorized as energy ($E$), force ($F$), or virial ($V$), with $p\in\{E,F,V\}$.
+For the specific domain of Molecules, the test sets are indexed as $i\in\{\text{MD22, ANI-1x, AIMD-Chig}\}$.
+To facilitate a fair comparison, the error metric is truncated and normalized against the error metric of a baseline model (dummy model) as follows:
+$\hat{M}^{m}_{k,p,i}=\min\left(\frac{M^{m}_{k,p,i}}{M^{\mathrm{dummy}}_{k,p,i}},1\right)$ (1)
 This baseline model predicts energy based solely on the chemical formula, disregarding any structural details, thereby providing a reference point for evaluating the improvement offered by more sophisticated models.
@@
 In this scenario, the prediction weights are adjusted to $w_{E}=w_{F}=0.45$ and $w_{V}=0.1$.
-The resulting error metric is referred to as $\bar{M}^{m}_{\mathrm{FF}}$.
+The resulting error metric after averaging over all domains is referred to as $\bar{M}^{m}_{\mathrm{FF}}$.
 For the domain-specific property calculation tasks, we adopt the MAE as the primary error metric.
-In the Inorganic Materials domain, the MDR phonon benchmark predicts maximum phonon frequency, entropy, free energy, and heat capacity at constant volume, with each prediction type assigned a weight of 0.25.
-In the Small Moleculesdomain, the TorsionNet 500 benchmark predicts the torsion profile energy, torsional barrier height, and the number of molecules for which the model’s prediction of the torsional barrier height has an error exceeding 1 kcal/mol.
-Each prediction type in this domain is assigned a weight of $\frac{1}{3}$.
-The resulting error metric is denoted as $\bar{M}^{m}_{\mathrm{PC}}$.
-IV.4 Applicability metrics
+In the Inorganic Materials domain, the MDR phonon benchmark predicts maximum phonon frequency, entropy, free energy, and heat capacity at constant volume, while the elasticity benchmark evaluates the shear and bulk moduli.
+Each prediction type is assigned an equal weight of 1/6.
+In the Molecules domain, the TorsionNet500 benchmark evaluates the torsion profile energy, torsional barrier height, and the number of molecules for which the predicted torsional barrier height error exceeds 1 kcal/mol.
+The Wiggle150 benchmark assesses the relative conformer energy profile.
+Each prediction type in this domain is assigned a weight of 1/4.
+In the Catalysis domain, the OC20NEB-OOD benchmark evaluates the energy barrier, reaction energy change (delta energy), and the percentage of reactions with predicted energy barrier errors exceeding 0.1 eV for three reaction types: transfer, dissociation, and desorption.
+Each prediction type in this domain is assigned a weight of 1/5.
+The resulting error metric after averaging over all domains is denoted as $\bar{M}^{m}_{\mathrm{PC}}$.
+IV.5 Applicability metrics
 The applicability metrics incorporate both efficiency and stability and are computed differently from the generalizability metrics due to the absence of a dummy model baseline.
-To evaluate efficiency, we define an efficiency metric, $M_{E}^{m}$ , by normalizing the average inference time (with unit $\mathrm{\mu s/atom}$), $\bar{\eta}^{m}$, of a given LAM measured over 1600 configurations with respect to an artificial reference value, thereby rescaling it to a range between zero and positive infinity.
+To evaluate efficiency, we define an efficiency metric, $M_{E}^{m}$ , by normalizing the average inference time (with unit $\mathrm{\mu s/atom}$), $\bar{\eta}^{m}$, of a given LAM measured over 900 configurations with respect to an artificial reference value, thereby rescaling it to a range between zero and positive infinity.
 A larger value indicates the higher efficiency.
-$M_{\mathrm{E}}^{m}=\frac{\eta^{0}}{\bar{\eta}^{m}},\quad\eta^{0}=100\ \mathrm{\mu s/atom},\quad\bar{\eta}^{m}=\frac{1}{1600}\sum_{i}^{1600}\eta_{i}^{m},$ (5)
+$M_{\mathrm{E}}^{m}=\frac{\eta^{0}}{\bar{\eta}^{m}},\quad\eta^{0}=100\ \mathrm{\mu s/atom},\quad\bar{\eta}^{m}=\frac{1}{900}\sum_{i}^{900}\eta_{i}^{m},$ (5)
 where $\eta_{i}^{m}$ is the inference time of configuration $i$ for model $m$.
@@
 $M^{m}_{\mathrm{IS}}=\frac{1}{9}\sum_{i=1}^{9}M^{m}_{\mathrm{IS},i}$ (7)
-This result is bounded within the range $[0,+\infty)$, where a lower value signifies greater stability.
+This result is bounded within the range $0,+\infty)$, where a lower value signifies greater stability.
+Author contributions
+A.P., C.C., L.Z., and H.W.
+conceived the idea of this work and developed the benchmark workflow.
+M.G., D.Z., and C.Z.
+collected and analyzed the results across the benchmark tasks.
+W.J.
+and Y.W.
+performed DFT calculations for dataset relabeling.
+A.L.
+provided guidance on the adaptation of the phonon benchmark.
+C.W.
+and W.E.
+provided guidance on the workflow design.
+All authors contributed to the discussions and revised the manuscript.
 Acknowledgments
@@
 Table S-1: Hyperparameters for DPA-2.4-7M model
-Module Hyperparameter DPA-2.4-7M Repinit nsel 120 neuron 3$\times$64 three_body_neuron 3$\times$32 Repformer nsel 40 nlayers 6 g1_dim 384 g2_dim 96 Activation tanh Training batch size “auto:256” nGPUs 120 $\times$ A800-40GB start lr 1e-3 stop lr 1e-5 precision float32 steps trained 8M
-Table S-2: Summary of datasets including dataset name, frame numbers, data cleaning, training average atoms, and DFT level.
-Dataset Name Frames Train Avg Atoms Data Cleaning Level of Theory
-Alex2D[65] trn: 1,223,831 val: 135,810 9.9 Exclude frames with energies $>$ 0 eV/atom, maximum absolute force $>$ 5 eV/Å, maximum absolute virial/atom $>$ 8 eV/atom; Trajectory first/last 3 frames + every 10 frame PBE/PAW, 520 eV, 0.4 Å-1
+[…]
+Table S-2: Summary of datasets, including the dataset name, number of frames, average number of atoms, data cleaning procedures, DFT level, and dataset weight used in training.
+Dataset Name # Frames Avg. # atoms Data Cleaning Level of Theory Weight
+Alex2D[77] trn: 1,223,831 val: 135,810 9.9 Exclude frames with energies >> 0 eV/atom, maximum absolute force >> 5 eV/Å, maximum absolute virial/atom >> 8 eV/atom; Trajectory first/last 3 frames + every 10 frame PBE/PAW, 520 eV, 0.4 Å-1 1.83
 MPTraj[13] trn: 1,401,956 val: 110,918 31.35 According to the distribution difference between MACE-MP-0 prediction and label, the union set of top 0.1% with the largest uncorrected energy difference and the top 0.1% with the largest force difference were excluded, resulting in a total of 3123 frames excluded.
-PBE (+U)/PAW, 520 eV, 0.04 Å-1
-OC20M[66] trn: 20,000,000 val: 999,866 73.1 - rPBE/PAW, 350eV
-OC22[67] trn: 8,194,770 val: 394,727 79.8 Exclude frames with energies $>$ 0eV/atom, maximum absolute force $>$ 10eV/Å PBE (+U)/PAW, 350 eV, 0.157 Å-1
-ODAC23[68] trn: 2,682,332 val: 63,623 203 Exclude frames with energy/atom $>$ 0.5 eV/atom or $<$ -0.2 eV/atom, and maximum absolute force $>$ 25 eV/Å; Trajectory first/last 3 frames + every 20th frame PBE-D3/PAW, 600 eV, gamma
-OMAT24[50] trn: 100,568,820 val: 1,074,647 18.6 Exclude frames with energies $>$ 0 eV/atom, $<$ -25eV/atom, maximum absolute force $>$ 50eV/Å PBE (+U)/PAW
-SPICE2[69] trn: 1,621,168 val: 180,089 35.6 Exclude frames with energies $<$ -10000 eV/atom, maximum absolute force $>$ 15eV/Å $\omega$B97M-D3(BJ)/def2-TZVPPD
-Transition1x[70] trn: 7,632,328 val: 967,454 13.91 - $\omega$B97X/6-31G(d)
-UniPero[71] trn: 14,487 val: 1,357 53.64 - PBEsol/LCAO, 1360 eV, 0.189 Å-1
-Dai2023Alloy[72] trn: 71,482 val: 1,240 20.99 - PBE/Norm-conserving, 1360 eV, 0.094 Å-1
-Zhang2023Cathode[73] trn: 88,692 val: 9,695 46.28 - PBE (+U)/PAW, 520 eV, 0.25 Å-1
-Gong2023Cluster[74] trn: 143,418 val: 15,331 23.67 - PBE-D3/TZV2P, 400-1000 Ry
-Li2025APEX[75] trn: 24,097 val: 100 24.03 - PBE/NC, 1360 eV, 0.15 Å-1
-Shi2024Electrolyte[76] trn: 65,393 val: 3,438 192.69 - PBE-D3, 800 Ry
-Shi2024SSE[77] trn: 125,083 val: 6,587 48.33 - PBE-sol/LCAO, 1360 eV, 0.28 Å-1
-Yang2023ab[78] trn: 1,379,956 val: 24,257 33.47 - $\omega$B97X-D/6-31G**
-Li2025General[79] trn: 14,024,587 val: 1,558,260 17.8 Exclude maximum absolute force $>$ 20eV/Å GFN2-xTB
-Huang2021Deep-PBE[80] trn: 17,582 val: 886 218.59 - PBE/PAW, 650 eV, 0.26 Å-1
-Liu2024Machine[81] trn: 215,481 val: 23,343 70.36 - PBE/LCAO, 1360 eV, 0.151 Å-1
-Zhang2021Phase[82] trn: 46,077 val: 2,342 172.3 - SCAN/PAW, 1500 eV, 0.5 Å-1
-Jiang2021Accurate[83] trn: 138,194 val: 3,965 23.16 - PBE/PAW, 650 eV, 0.1 Å-1
-Chen2023Modeling[84] trn: 6,449 val: 276 16.94 - SCAN/PAW, 650 eV, 0.1 Å-1
-Unke2019PhysNet[85] trn: 2,594,609 val: 136,571 21.38 - revPBE-D3(BJ)/def2-TZVP
-Wen2021Specialising[86] trn: 10,054 val: 474 20.02 - PBE/PAW, 650 eV, 0.1 Å-1
-Wang2022Classical[87] trn: 14,935 val: 738 24.47 - PBE/PAW, 650 eV, 0.1 Å-1
-Wang2022Tungsten[88] trn: 42,297 val: 2,100 24.91 - PBE/PAW, 600 eV, 0.16 Å-1
-Wu2021Deep[89] trn: 27,660 val: 917 96 - PBE/PAW, 600 eV, 2$\times$2$\times$2 kpt
-Huang2021Deep-PBEsol[80] trn: 7,502 val: 384 160.8 - PBE-sol/PAW, 650 eV, 0.26 Å-1
-Wang2021Generalizable[90] trn: 64,239 val: 2,256 21.85 - PBE-D3/PAW, 650 eV, 0.1 Å-1
-Wu2021Accurate[91] trn: 11,621 val: 568 21.73 - PBE/PAW, 700 eV
-Tuo2023Hybrid[92] trn: 48,078 val: 2,530 45.11 - PBE-D3(BJ)/PAW, 500 eV, 0.16 Å-1
+PBE (+U)/PAW, 520 eV, 0.04 Å-1 4.66
+OC20M[65] trn: 20,000,000 val: 999,866 73.1 - rPBE/PAW, 350eV 11.65
+OC22[78] trn: 8,194,770 val: 394,727 79.8 Exclude frames with energies >> 0eV/atom, maximum absolute force >> 10eV/Å PBE (+U)/PAW, 500 eV 8.14
+ODAC23[79] trn: 2,682,332 val: 63,623 203 Exclude frames with energy/atom >> 0.5 eV/atom or << -0.2 eV/atom, and maximum absolute force >> 25 eV/Å; Trajectory first/last 3 frames + every 20th frame PBE-D3/PAW, 600 eV, gamma 1.19
+OMAT24[72] trn: 100,568,820 val: 1,074,647 18.6 Exclude frames with energies >> 0 eV/atom, << -25eV/atom, maximum absolute force >> 50eV/Å PBE (+U)/PAW 22.16
+[…]
+[…]
+UniPero[82] trn: 14,487 val: 1,357 53.64 - PBEsol/LCAO, 1360 eV, 0.189 Å-1 0.77
+Dai2023Alloy[83] trn: 71,482 val: 1,240 20.99 - PBE/Norm-conserving (NC), 1360 eV, 0.094 Å-1 0.67
+Zhang2023Cathode[84] trn: 88,692 val: 9,695 46.28 - PBE (+U)/PAW, 520 eV, 0.25 Å-1 1.64
+Gong2023Cluster[85] trn: 143,418 val: 15,331 23.67 - PBE-D3/TZV2P, 400-1000 Ry 1.07
+Li2025APEX[86] trn: 24,097 val: 100 24.03 - PBE/NC, 1360 eV, 0.15 Å-1 0.42
+Shi2024Electrolyte[87] trn: 65,393 val: 3,438 192.69 - PBE-D3, 800 Ry 5.85
+Shi2024SSE[88] trn: 125,083 val: 6,587 48.33 - PBE-sol/LCAO, 1360 eV, 0.28 Å-1 2.03
+[…]
+Li2025General[90] trn: 14,024,587 val: 1,558,260 17.8 Exclude maximum absolute force >> 20eV/Å GFN2-xTB 7.95
+Huang2021Deep-PBE[91] trn: 17,582 val: 886 218.59 - PBE/PAW, 650 eV, 0.26 Å-1 1.72
+Liu2024Machine[92] trn: 215,481 val: 23,343 70.36 - PBE/LCAO, 1360 eV, 0.151 Å-1 3.88
+Zhang2021Phase[93] trn: 46,077 val: 2,342 172.3 - SCAN/PAW, 1500 eV, 0.5 Å-1 4.39
+Jiang2021Accurate[94] trn: 138,194 val: 3,965 23.16 - PBE/PAW, 650 eV, 0.1 Å-1 0.51
+Chen2023Modeling[95] trn: 6,449 val: 276 16.94 - SCAN/PAW, 650 eV, 0.1 Å-1 0.03
+Unke2019PhysNet[96] trn: 2,594,609 val: 136,571 21.38 - revPBE-D3(BJ)/def2-TZVP 4.09
+Wen2021Specialising[97] trn: 10,054 val: 474 20.02 - PBE/PAW, 650 eV, 0.1 Å-1 0.05
+Wang2022Classical[98] trn: 14,935 val: 738 24.47 - PBE/PAW, 650 eV, 0.1 Å-1 0.07
+Wang2022Tungsten[99] trn: 42,297 val: 2,100 24.91 - PBE/PAW, 600 eV, 0.16 Å-1 0.12
+Wu2021Deep[100] trn: 27,660 val: 917 96 - PBE/PAW, 600 eV, $\times2$$\times$2 kpt 0.95
+Huang2021Deep-PBEsol[91] trn: 7,502 val: 384 160.8 - PBE-sol/PAW, 650 eV, 0.26 Å-1 0.83
+Wang2021Generalizable[101] trn: 64,239 val: 2,256 21.85 - PBE-D3/PAW, 650 eV, 0.1 Å-1 0.26
+Wu2021Accurate[102] trn: 11,621 val: 568 21.73 - PBE/PAW, 700 eV 0.14
+Tuo2023Hybrid[103] trn: 48,078 val: 2,530 45.11 - PBE-D3(BJ)/PAW, 500 eV, 0.16 Å-1 0.59
 S-2 Original Error Metrics for Generalizability Domain Specific Tasks
-Table S-3: Phonon related property prediction on MDR phonon benchmark.
+Table S-3: Phonon related property prediction on the MDR phonon benchmark.
 Model MAE $\omega_{\text{max}}$ (K) MAE $S$ (J/K/mol) MAE $F$ (kJ/mol) MAE $C_{V}$ (J/K/mol)
+DPA-3.1-3M-DomainMatch 32.4 24.4 7.0 9.1
+DPA-3.1-3M 38.4 33.4 9.9 10.9
 DPA-2.4-7M 58.9 45.6 18.1 12.8
@@
 Orb-v2 308.0 446.5 184.3 58.5
+Orb-v3 72.1 31.5 7.3 13.6
 SevenNet-l3i5 25.6 25.9 9.6 4.9
@@
 dummy 1188.3 764.8 125.1 547.4
-Table S-4: Torsional MAE and MAEB errors between LAM predictions and its reference DFT labels on the TorsionNet-500 Benchmark.
+Table S-4: Benchmark results on elasticity related property prediction.
+Model MAE GVRH MAE KVRH
+DPA-3.1-3M-DomainMatch 9.83 8.68
+DPA-3.1-3M 10.77 10.13
+DPA-2.4-7M 17.76 16.46
+MACE-MP-0 26.20 11.01
+MACE-MPA-0 10.27 15.03
+Orb-v2 66.07 44.08
+Orb-v3 9.75 7.58
+SevenNet-l3i5 19.42 9.93
+SevenNet-MF-ompa 9.54 9.46
+MatterSim-v1-5M 12.75 14.95
+GRACE-2L-OAM 9.14 7.46
+dummy 67.54 136.26
+Table S-5: Torsional MAE and MAEB errors between LAM predictions and its reference DFT labels on the TorsionNet-500 Benchmark.
 Model MAE (kcal/mol) MAEB11footnotemark: 1 (kcal/mol) $\text{NABH}_{h}$22footnotemark: 2
-DPA-2.4-7M 0.727 1.220 267
-MACE-MP-0 1.653 2.425 354
-MACE-MPA-0 1.486 2.179 339
-Orb-v2 1.238 2.067 345
-SevenNet-l3i5 1.113 1.641 300
-SevenNet-MF-ompa 1.599 2.309 354
-MatterSim-v1-5M 1.348 2.240 345
-GRACE-2L-OAM 1.387 1.992 303
-dummy 2.501 5.877 494
+DPA-3.1-3M-DomainMatch 0.338 0.551 76
+DPA-3.1-3M 0.825 1.382 298
+DPA-2.4-7M 0.739 1.229 265
+MACE-MP-0 1.652 2.437 356
+MACE-MPA-0 1.468 2.150 339
+Orb-v2 1.248 2.109 346
+Orb-v3 0.988 1.421 268
+SevenNet-l3i5 1.127 1.641 308
+SevenNet-MF-ompa 1.619 2.338 360
+MatterSim-v1-5M 1.360 2.234 342
+GRACE-2L-OAM 1.394 1.985 309
+dummy 2.494 5.818 493
 11footnotemark: 1
@@
 22footnotemark: 2The number of molecules (total: $N_{\text{mols}}$ = 500) for which the model prediction of torsional barrier height has an error of more than 1 kcal/mol.
-S-3 Downsampling Details for OOD Datasets
-We applied a variety of down-sampling strategies across eleven benchmark datasets to reduce computational cost while preserving representativeness.
-The datasets employed in this study underwent various downsampling procedures to ensure manageable sizes while maintaining representative distributions.
-For the ANI-1x dataset, 10% of frames were randomly sampled from the validation set using a fixed random seed 42.
-The Lopanitsyna2023Modeling_B dataset was used in its entirety without downsampling, while its surface counterpart Lopanitsyna2023Modeling_A utilized the validation split provided by the original authors.
-The combined Dai2024Deep dataset was reduced to 5% of its original size through random sampling with seed 0.
-For the MD22 dataset, 5% of frames were randomly sampled from systems of short peptides (Ac-Ala3-NHMe), docosahexaenoic acid, stachyose, and DNA base pairs (AT-AT and AT-AT-CG-CG), with random seed set to 0 and units fixed.
-The Gasteiger2020Fast dataset retained its original test split without any downsampling.
-The Guan2022Benchmark dataset was 5% subset randomly sampled (seed 0) from the pre-cleaned data which excluding frames with energy $>-0.5~{}\mathrm{eV/atom}$ and maximum force $>30~{}\mathrm{eV/\text{\AA}}$.
-The Zhang2019Bridging dataset excluded frames with energy $<0.53~{}\mathrm{eV/atom}$ and maximum forces $>5~{}\mathrm{eV/\text{\AA}}$.
-The Batzner2022equivariant dataset was downsampled to 5% (seed 0) after removing frames with energy $<-50~{}\mathrm{eV/atom}$ and maximum forces $>10~{}\mathrm{eV/\text{\AA}}$.
-The AIMD-Chig dataset underwent downsampling at 0.5% (seed 0).
-The Zhang2024Active dataset was randomly downsampled to 5% (seed 0) after removing frames with energy $>-2.5~{}\mathrm{eV/atom}$ and maximum forces $>25~{}\mathrm{eV/\text{\AA}}$.
-The Villanueva2024Water dataset was used in full after removing frames with energy ($<-6~{}\mathrm{eV/atom}$) and force ($<2~{}\mathrm{eV/\text{\AA}}$).
-The Subalex_9k dataset was created by randomly sampling 9157 frames (seed 42) from pre-cleaned data that excluded frames with energy $<-25~{}\mathrm{eV/atom}$ and maximum forces $>3~{}\mathrm{eV/\text{\AA}}$ from sAlex.
-The WBM_25k dataset was downsampled to 25696 frames (seed 42).
-The Sours2023Applications dataset was reduced to 5% (seed 0) after removing frames with energy $>-5~{}\mathrm{eV/atom}$.
-The Torres2019Analysis (Ca_Battery) was used in all GGA-labeled dataset with no virials.
-The Vandermause2022Active was used in its full set of DFT labels.
-S-4 Structures for Conservativeness Benchmark
+Table S-6: Prediction of conformer relative energies on the Wiggle150 benchmark.
+Model MAE (kcal/mol)
+DPA-3.1-3M-DomainMatch 1.72
+DPA-3.1-3M 5.67
+DPA-2.4-7M 14.84
+MACE-MP-0 26.60
+MACE-MPA-0 14.91
+Orb-v2 6.46
+Orb-v3 11.92
+SevenNet-l3i5 13.88
+SevenNet-MF-ompa 10.97
+MatterSim-v1-5M 10.73
+GRACE-2L-OAM 12.14
+dummy 105.06
+Table S-7: Energy barrier prediction on the OC20NEB-OOD Benchmark.11footnotemark: 1
+Model MAEEa22footnotemark: 2 (eV) MAEdE33footnotemark: 3 (eV) $\phi_{\text{Transfer}}$44footnotemark: 4 $\phi_{\text{Dissociation}}$44footnotemark: 4 $\phi_{\text{Desorption}}$44footnotemark: 4
+DPA-3.1-3M-DomainMatch 1.172 0.158 65.1 69.0 59.8
+DPA-3.1-3M 1.203 0.234 66.9 63.9 81.1
+DPA-2.4-7M 1.271 0.306 68.6 77.9 69.3
+MACE-MP-0 1.468 0.580 76.6 84.8 90.6
+MACE-MPA-0 1.459 0.566 72.6 84.8 94.5
+Orb-v2 2.682 1.729 66.3 74.1 78.7
+Orb-v3 2.298 1.470 61.7 72.2 87.4
+SevenNet-l3i5 1.423 0.546 77.7 76.0 87.4
+SevenNet-MF-ompa 2.070 1.278 66.9 68.4 92.9
+MatterSim-v1-5M 2.686 2.018 76.0 84.8 83.5
+GRACE-2L-OAM 1.583 0.704 65.1 72.2 90.6
+dummy 2.360 0.950 95.4 99.4 95.3
+11footnotemark: 1
+Only 460 out-of-distribution trajectories were used in this task.
+22footnotemark: 2The mean absolute error of the reaction barrier.
+33footnotemark: 3The mean absolute error of the reaction energy.
+44footnotemark: 4Percentage of reactions with an energy barrier error exceeding 0.1 eV.
+The analysis considers three reaction types: transfer, dissociation, and desorption.
+S-3 Structures for Conservativeness Benchmark
 Figure S-1: Structures used for the Conservativeness Benchmark. (a) H2Al32Cr48Mn16N2O; (b) Cs8N2; (c) Gd2Ni2Si4; (d) NdPd3; (e) BaNi2O8V2; (f) BaNiO5Tm2; (g) CH3N5S; (h) C3H5N2; (i) C4H7NO.
-S-5 Inference Efficiency Convergence Test
+S-4 Inference Efficiency Convergence Test
 Figure S-2: The structures in unit cell for convergence test in efficiency benchmark.
@@
 Figure S-3: The convergence test of efficiency benchmark with respect to atom numbers.
-S-6 The Coordination Number Analysis of Configurations in Efficiency Benchmarks
+S-5 The Coordination Number Analysis of Configurations in Efficiency Benchmarks
 Figure S-4: Effect of the number of neighbor atoms on inference time, (a) correlation between inference time and vacuum thickness in a bilayer $\text{Na}_{8}$ 2D structure.
@@
 The number of neighbor atoms is estimated using the cutoff radius specific to each model.
-S-7 Property Fine-tuning Batch Size Effect
-As mentioned earlier, prediction accuracy is strongly correlated with the training batch size due to the limited size of the fine-tuning dataset.
-The default workflow utilizes NVIDIA L4 GPUs for cost management purposes, which results in a notable decline in performance.
-However, the relative performance trends among models remain consistent.
-For the MP $E_{\text{form}}$ and MP Gap tasks, the reported accuracy may not reflect the fully converged results due to insufficient training epochs.
-Table S-5: Comparison of property fine-tuning accuracy on NVIDIA L4-24GB and A800-40GB GPU.
-For MP $E_{\text{form}}$ and MP Gap, the batch size on the L4 was set to a quarter of that used on the A100, while for all other tasks, it was set to half.
-Tasks (Units) DPA-2.4-7M L4 Finetune DPA-2.4-7M A800 Finetune Estimated Epochs11footnotemark: 1 L4 |||| A800 MP $E_{\text{form}}$ (meV/atom) 31.1 24.1 80 |||| 320 MP Gap (eV) 0.285 0.206 100 |||| 400 JDFT2D (meV/atom) 46.93 34.09 28500 |||| 57000 Phonons (cm-1) 41.10 34.00 14000 |||| 28000 Dielectric (unitless) 0.431 0.300 3000 |||| 6000 Log KVRH ($\log_{10}\text{GPA}$) 0.062 0.052 1300 |||| 2600 Log GVRH ($\log_{10}\text{GPA}$) 0.079 0.068 1300 |||| 2600 Perovskites (eV/unitcell) 0.061 0.032 1350 |||| 2700
-11footnotemark: 1
-The DPA models employ a dynamic batch size to optimize training efficiency, making it difficult to precisely determine the number of training epochs.
-S-8 Stability Test Results
-Table S-6: NVE molecular dynamics simulations over nine atomic systems.
+S-6 Stability Test Results
+Table S-8: NVE molecular dynamics simulations over nine atomic systems.
 Model Energy Drift (meV/atom/ps) Success Rate Direct Force Prediction
+DPA-3.1-3M 0.008 0.89 No
 DPA-2.4-7M 0.018 1 No
@@
 Orb-v2 222.8 1 Yes
+Orb-v3 0.005 1 No
 SevenNet-l3i5 0.012 1 No